### v1

# Review, Refine, Repeat: Understanding Iterative Decoding of AI Agents with Dynamic Evaluation and Selection

Abstract
While AI agents have shown remarkable performance at various tasks, they still struggle with complex multi-modal applications, structured generation and strategic planning. Improvements via standard fine-tuning is often impractical, as solving agentic tasks usually relies on black box API access without control over model parameters. Inference-time methods such as Best-of-N (BON) sampling offer a simple yet effective alternative to improve performance. However, BON lacks iterative feedback integration mechanism. Hence, we propose Iterative Agent Decoding (IAD) which combines iterative refinement with dynamic candidate evaluation and selection guided by a verifier. IAD differs in how feedback is designed and integrated, specifically optimized to extract maximal signal from reward scores. We conduct a detailed comparison of baselines across key metrics on Sketch2Code, Text2SQL, and Webshop where IAD consistently outperforms baselines, achieving 3–6% absolute gains on Sketch2Code and Text2SQL (with and without LLM judges) and 8–10% gains on Webshop across multiple metrics. To better understand the source of IAD’s gains, we perform controlled experiments to disentangle the effect of adaptive feedback from stochastic sampling, and find that IAD’s improvements are primarily driven by verifier-guided refinement, not merely sampling diversity. We also show that both IAD and BON exhibit inference-time scaling with increased compute when guided by an optimal verifier. Our analysis highlights the critical role of verifier quality in effective inference-time optimization and examines the impact of noisy and sparse rewards on scaling behavior. Together, these findings offer key insights into the trade-offs and principles of effective inference-time optimization.
1 Introduction
AI agents have demonstrated remarkable capabilities across a wide range of tasks, including natural language understanding, code generation, and mathematical reasoning (Fu et al., 2024; Liu et al., 2024; Jiang et al., 2024; Patel et al., 2024; Huang et al., 2024; Liu et al., 2023). However, state-of-the-art AI agents still face challenges when handling tasks that require understanding multimodal inputs and reasoning through structured output generation (Costarelli et al., 2024; Kinniment et al., 2024; Liu et al., 2023; Verma et al., 2024; Valmeekam et al., 2024). For example, consider agentic tasks such as Sketch2Code (Li et al., 2024b), Text2SQL (Li et al., 2024a), and the web-navigation task Webshop (Yao et al., 2023). Sketch2Code requires an AI agent to generate structured HTML from rough sketches, often resulting in incorrect generations. Text2SQL involves translating natural language queries into executable SQL statements, demanding precise reasoning while avoiding semantic and logical errors. Meanwhile, Webshop is a sequential decision-making task that evaluates an agent’s ability to navigate an e-commerce platform effectively. Across these tasks, AI agents consistently underperform compared to human experts, underscoring the limitations of current agentic models in handling complex, structured outputs and sequential reasoning (Yao et al., 2023; Liu et al., 2024; Li et al., 2024b) where agents achieve merely 20-30% performance in task specific metrics.
Figure 1: Qualitative (Left) and Quantitative (Right) illustration of the performance benefit of IAD (Ours) over Single-turn response generations using Gemini-1.5 (Base) Text2SQL task (Li et al., 2024a). IAD improves performance by correctly handling query logic and joins, improving the accuracy over baseline and BON (Best-of-N).
A key challenge in improving agentic AI systems is their ability to handle structured reasoning and multimodal tasks effectively. One promising approach to addressing this challenge is fine-tuning a new model through alignment (Ouyang et al., 2022a). A key alignment technique, reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022a, b; Bai et al., 2022) has significantly enhanced the capabilities of generative models, it is not directly applicable to AI agents due to their inherently black box nature. Most agentic systems and frameworks (Liu et al., 2024; Fu et al., 2024) operate in environments where internal model updates are infeasible, making fine-tuning-based alignment techniques inapplicable. An alternative approach is inference-time alignment, such as controlled decoding (Khanov et al., 2024; Deng and Raffel, 2023; Mudgal et al., 2024; Han et al., 2024; Chakraborty et al., 2024; Yang and Klein, 2021), which can align generative models without fine-tuning. However, these methods still rely on access to token-level logits from the reference model to steer responses, an assumption that does not hold for black box agentic systems. These limitations highlight the need for an inference-time alignment technique that is (i) compatible with various underlying models (without requiring model-specific adaptations), (ii) easy to implement (requires no modification of the underlying models), and (iii) commercial API-friendly (Applicable to closed-source models that are only accessible via API. This leads us to an important question:
How can we improve the inference alignment of AI Agents with black box access for complex and structured tasks?
A simple solution to the above question is Best-of-N (BON) sampling and its variants (Nakano et al., 2021; Stiennon et al., 2020; Beirami et al., 2024), as it is black box compatible and logit-free, making it a natural choice for inference-time alignment. BON selects the best response from multiple generated candidates using a verifier, improving response quality in structured tasks such as HTML generation and SQL query synthesis. BON is even almost optimal in terms of quality (Gui et al., 2025; Yang et al., 2024b; Beirami et al., 2024; Amini et al., 2024). However, BON may be ineffcient requiring a large N to achieve reasonable performance. To address these limitations, we propose Iterative Agent Decoding (IAD)—an iterative approach that refines responses through verifier-guided feedback. IAD differs fundamentally in the design and integration of its feedback mechanism, specifically optimized to extract and utilize rich signals from the reward score thus demonstrating the effectiveness of adaptive feedback in complex tasks. We then conduct a detailed comparison of baselines including parallel sampling (BON) and iterative decoding (IAD) across several key metrics relevant to inference-time scaling, yielding key insights into the trade-offs and principles of effective inference-time optimization. We summarize our contributions and findings as follows.
Figure 2: Demonstrates the high-level schematic of our proposed approach - Iterative Decoding of AI Agents
1. An iterative decoding framework (IAD): We propose Iterative Agent Decoding (IAD), an efficient iterative framework where responses are refined through verifier-guided feedback, enabling improved performance in black box structured generation tasks. IAD is designed to extract maximal information from available feedback, including sparse or noisy scores.
2. Empirical evaluation across agentic tasks: We evaluate IAD on three complex structured output tasks, demonstrating consistent improvements over strong baselines, while maintaining efficiency. IAD minimizes reliance on crafted exemplars and leverages feedback-driven reconditioning, making it task-agnostic and model-agnostic.
3. Disentangling feedback gains in IAD from sampling: We conduct controlled experiment to isolate the effect of adaptive feedback in IAD from sampling diversity under low-temperature sampling, where IAD continues to improve, while BON performance stagnates, confirming that IAD’s gains are driven by adaptive, feedback-guided refinement, not sampling diversity alone.
4. Inference scaling behavior: We highlight the inference-time scaling of performance improvements where we show that both IAD and BON improve with increased test-time compute when guided by an optimal verifier. However, IAD benefits more significantly from high-quality feedback, leading to greater performance gains. Additionally, as model capability increases, BON performance improves, narrowing the gap with IAD.
5. Importance of verifier quality for inference scaling: We show that high-quality verifier feedback is critical for inference scaling. Self-evaluation or sub-optimal verifiers often fail in complex agentic scenarios, whereas optimal verifier-guided alignment enables consistent refinement.
6. Robustness under sparse and noisy feedback: To further illustrate the impact of verifier quality, we conduct controlled experiments by sparsifying and adding noise to the verifier signal. Results show that IAD remains robust under mild sparsity and noise, but performance for both IAD and BON degrades under extreme conditions, due to the lack of reliable feedback.
Remarks: Several multi-turn/multi-attempt iterative approaches focus on train-time methods (Qu et al., 2024; Zhou et al., 2024; Shani et al., 2024), making them unsuitable for black box inference. Inference-time methods (Madaan et al., 2023; Chao et al., 2024; Mehrabi et al., 2024; Muennighoff et al., 2025) have explored iterative refinement in specific contexts like jailbreak, math, and QA, but primarily for simpler tasks rather than complex agentic settings requiring structured generations like HTML and SQL. Specifically, Chao et al. (2024); Mehrabi et al. (2024) iteratively refine prompts for adversarial attacks. Meanwhile, Madaan et al. (2023) and its variants emphasize on LLM self-evaluation rather than focusing on the design choices for iterative refinement based approach for structured output generation.
2 AI Agent Alignment
In this section, we formalize the problem of inference-time alignment for AI agents operating in black box settings, particularly in complex structured-generation tasks such as HTML and SQL synthesis in multi-modal environments like Sketch2Code, Text2SQL, and Webshop. Given a task distribution $\mathcal{T}\in\Delta$, we define the optimal but inaccessible agent policy $\pi^{*}(\cdot|x)$, which generates the ideal structured response $y^{*}\sim\pi^{*}(\cdot|x)$ for a given input prompt $x$. However, in the black box setting, we only have access to a reference policy $\pi_{0}(\cdot|x)$, which may lack some capabilities needed to complete the task.
Figure 3: Sketch2code: This figure provides a comparison of IAD (ours) against Best-of-N sampling (SoTA) and single-turn generation with Gemini-1.5-Pro with temperature =0.6 w.r.t metrics - (a) Layout Similarity (b) TextIoU (c) ImageIoU across varying the number of generations (N). Figure demonstrates IAD outperforms BON consistently across N, but as N increases the gap reduces. Note that (b) and (c) are not directly optimized for this plot and test the generalization performance. This demonstrates the consistency in performance across models and metrics
The core challenge is to optimize inference-time decisions such that the generated response from $\pi_{0}(\cdot|x)$ is better aligned with the outputs of $\pi^{*}(\cdot|x)$. To formally quantify this alignment gap, we define the inference-time misalignment as an $f$-divergence between the optimal and reference policies as $d_{f}(\pi^{*}(\cdot|x),\pi_{0}(\cdot|x))$, where $d_{f}(\cdot,\cdot)$ measures how far the black box agent deviates from the ideal structured output distribution. We explore methods to reduce this divergence using iterative inference-time strategies, which we detail in the following sections. We assume access to a verifier function $R(x,y)$, which evaluates the quality of a response $y$ given the prompt $x$ (further details in Appendix).
Task specific details: In Sketch2code, the reference policy $\pi_{0}(\cdot|x)$ generates structured HTML $y$ given the prompt $x=[s,c]$ with wireframe sketch $s$, noisy text prompt $c$. The verifier $R(x,y)$ evaluates alignment with the target layout, textual coherence, and semantic correctness, ensuring that the generated HTML accurately reflects the sketch’s structure. In Text2SQL, $\pi_{0}(\cdot|x)$ maps a natural language query $c$ to an SQL query $y$ conditioned on the database schema $x$. The verifier $R(x,y)$ assesses syntactic validity and execution correctness, identifying errors in joins, logical consistency, and query structure. Finally, in WebShop, the agent’s actions are conditioned on the full interaction history up to the current time step, i.e $x_{t}=[x,y_{<t}]$, where $x$ is the initial user query and $y_{<t}$ represents all previously generated product page interactions. Based on this history, the reference policy $\pi_{0}(\cdot|x_{t})$ generates the next action $a_{t}$ which involves searching or clicking and the corresponding product description.
Model Layout. Txt IoU Img IoU
Single-Turn Approaches
InternVL2-8b 4.01 4.89 1.41
Llava-1.6-8b 8.01 9.26 1.95
Claude-3-Sonnet 14.22 15.85 6.62
GPT-4o-Mini 16.29 20.84 0.72
Claude-3-Opus 17.11 18.09 8.32
Claude-3-Haiku 17.52 20.60 2.72
Gemini-1.5-Flash 17.9 17.50 10.77
Gemini-1.5-Pro 18.25 18.20 12.69
GPT-4o 19.20 17.12 16.19
Gemini-1.5-Flash (CoT) 19.84 19.13 10.02
Claude-3.5-Sonnet 22.26 25.33 9.21
Multi-Turn Approaches (Gemini-1.5-Flash)
Sk2code (N=2)∗∗ 19.41 20.45 11.81
Self-Refine (N=2) 19.51 19.35 10.71
BON (N=2) 21.45 20.1 13.5
IAD (N=2) 24.78 23.01 15.29
IAD-fb (N=2, K=2)∗∗ 24.66 23.4 14.69
Self-Refine (N=4) 19.97 19.11 11.74
Sk2code (N=4)∗∗ 20.41 21.46 12.67
BON (N=4) 24.02 22.59 15.91
IAD (N=4) 25.97 24.13 16.98
IAD-fb (N=4, K=4)∗∗ 26.61 24.62 17.36
Self-Refine (N=6) 19.89 18.91 11.61
Sk2code (N=6)∗∗ 21.43 21.53 13.78
BON (N=6) 25.75 22.91 17.67
IAD (N=6) 26.75 24.91 19.12
IAD-fb (N=6, K=6)∗∗ 27.95 24.99 19.01
Table 1: Sketch2Code: Performance comparison between single-turn and multi-response generation approaches. For each of the multi-response generation method Layout score acts as the reference metric (temperature =0.6). Table demonstrate that IAD (Ours) consistently outperform SoTA baseline by >3-4% margin (absolute). ∗∗ indicates those approaches leverage LLM-judge (oracle) feedback, although the LLM judges are similar but not exactly the same. Numbers taken from (Li et al., 2024b). N denotes the number of LLM calls for generating the HTML (>2000 tokens) and K represents the calls to LLM judge for getting feedback (<200 tokens). We have kept the judge-prompt similar to Sk2code (Li et al., 2024b)
2.1 Limitation of Prior approaches
In this section, we first provide a brief description of the baseline approaches and then discuss their pros and cons in this context.
Single-turn Approaches: In single-turn approaches, the response $y\sim\pi_{0}(\cdot|x)$ is directly generated from the reference agent policy, appending some instructions to accomplish the task. This method is straightforward, fast, and does not rely on a verifier, making it applicable even in verifier-agnostic settings. However, as evident from Figure 1, direct generation — even with models like Gemini-1.5-Pro, Gemini-1.5-Flash, GPT-4, and Claude 3.X — remains highly sub-optimal for complex tasks like Sketch2Code and Text2SQL, also highlighting the difficulty of these tasks. Thus in single-turn generation, the performance is limited by the quality of the reference policy $\pi_{0}(\cdot|x))$ where larger f-divergence indicates greater misalignment.
BON sampling: Best-of-N sampling improves upon single-turn generation by drawing $N$ i.i.d samples from the reference policy $\pi_{0}(\cdot|x)$ and selecting the highest-reward response based on the verifier $R(x,y)$. BON is simple, parallelizable, and computationally efficient and doesn’t rely on logits/model access thus applicable to black box agentic scenarios. It works even with scalar rewards and has been shown to achieve near-optimal trade-offs between win rate and KL divergence (Yang et al., 2024b; Beirami et al., 2024; Gui et al., 2025). Despite its advantages, BON remains limited by the quality of the reference policy lacks the ability to iteratively refine responses based on verifier feedback. For example: BON cannot incorporate targeted feedback, such as refining specific HTML structures in Sketch2Code or correcting systematic SQL errors in Text2SQL (further details in exp section)
Figure 4: Sketch2code: This figure provides a comparison of IAD with Best-of-N sampling (BON) and single-turn generation with Gemini-1.5-Flash with temperature =0.6 for the metrics - (a, d) Layout Similarity (b, e) TextIoU (c, f) ImageIoU across varying the number of generations (N). Top 3 rows, the optimization is done taking Text IOU as the verifier and the bottom 3 rows with Image IOU as the verifier. So, this also shows both the generalisability and performance improvement of IAD over baselines.
3 Proposed Methodology
To address the limitations of the existing approaches, as discussed in the previous section, we propose an iterative procedure for inference time alignment of AI agents in this section.
Our key insight: We begin by emphasizing that any iterative optimization approach generally follows three basic steps: (1) a generator to produce responses, (2) a verifier or reward function to evaluate them, and (3) a feedback mechanism to guide the next generation. This high-level loop is standard across any iterative frameworks, and novelty typically arises in how these components—particularly the feedback mechanism—are instantiated and integrated. Mathematically, this implies that if we can design a feedback mechanism over multiple iterations, denoted as $f(y_{1},y_{2},\dots,y_{T})$ , such that the such that the divergence $d_{f}(\pi^{*}(\cdot|x),\pi_{0}(\cdot|x,f(y_{1},y_{2},\dots,y_{T})))$ reduces, then we can say that feedback mechanism is effective. progressively decreases, then the feedback mechanism can be considered effective. In other words, the iterative process is successfully guiding the responses towards optimality. Motivated by this, we introduce our proposed approach: Iterative Agent Decoding (IAD).
Proposed approach:
Figure 5: Sketch2code: (a) Compares the performance of IAD at low temperature across varying number of generations to disentangle the effect of stochasticity and the improvement from iterative feedback based on IAD.(b) Comparison with BON at temperature = 0.1, which shows BON improvement ceases after 2 iterations Both (a, b) highlights the improvement due to adaptive refinement with verifier(c) Illustrates the performance of our proposed approach over 4-turns w.r.t Layout similarity score for 6 examples (temperature = 0.6).
We propose an iterative refinement approach for black box AI agents, where response generation is guided by a verifier to progressively improve outputs.
Step 1: At each step $t$, we sample a candidate response with the reference policy $\pi_{0}(\cdot|x)$ and iteratively refine its outputs.
$y_{t+1}\sim\pi_{0}(\cdot|x,\hat{y}_{t},s_{t},p_{t}),$ (1)
where $\hat{y}_{t}$ is the best response from previous iterations, and $p_{t}$ encodes prompt-based guiding instructions for ex: Surpass the best response, avoiding previous mistakes. This represents the more simplistic setting. Specifically in our algorithm, we condition on the best response $\hat{y}_{t}^{b}$ and worst response $\hat{y}_{t}^{w}$ over the iterations along with the basic instruction prompt to provide clear signals.
Step 2: Verifier-guided selection Among the generated response and the previous best response, we select the one maximizing the reward function:
$\hat{y}_{t+1}=\arg\max_{y\in(y_{t},\hat{y}_{t})}R(x,y)$ (2)
Our method also has the flexibility to incorporate critique-based feedback, such as LLM-judge, which identifies specific areas needing improvement. More specifically we keep a track of both the best response $\hat{y}_{t+1}^{b}=\arg\max_{y\in(y_{t},\hat{y}_{t}^{b})}R(x,y)$ and similarly the worst response $\hat{y}_{t+1}^{w}=\arg\min_{y\in(y_{t},\hat{y}_{t}^{w})}R(x,y)$. Thus this extends to
$y_{t+1}\sim\pi_{0}(\cdot|x,\hat{y}_{t}^{b},\hat{y}_{t}^{w},p_{t}fb_{t})$ (3)
where $fb_{t}$ highlights specific components (e.g., incorrect tags in HTML, invalid SQL joins), guiding refinement at a more granular level in an iterative fashion.
3. Acceptance Criterion : A new response $y_{t+1}$ is accepted if it provides an improvement over the previous best, $R(x,y_{t+1})-R(x,\hat{y}_{t})>0$. T
4. Iterative Refinement: The accepted response $\hat{y}_{t}$ updates the context for subsequent generations, progressively re-weighting the proposal distribution toward higher-quality outputs. By conditioning on $\hat{y}_{t}$, the sampling process is guided towards responses with higher rewards, reducing the performance gap via achieving higher rewards, as shown in experimental results.
Figure 6: Sketch2code: (a)Represents the performance of IAD with N = 2 w.r.t Layout score with varied sparsity NS (No Sparsity), LS (Low Sparsity) HS (High Sparsity) and ES (Extreme) (b) Performance comparison of IAD with BON with N = 2 at varied sparsity levels which shows the gains over BON reduces with high sparsity (c) Performance comparison of IAD with BON with N = 2 at varied noise levels of reward score which shows gains reduces over BON. However, IAD still improves over baselines under Sparse and Noisy feedback
Remark (Why it works?). The generation of better responses occurs through stochastic sampling in each iteration from the base model, conditioned on the best response so far (and the worst candidate), followed by a pairwise comparison from the verifier. This feedback mechanism helps guide the generator to sample responses with higher expected rewards over the iterations. Practically, we achieve this by incorporating prompts like “Improve upon the best response while avoiding mistakes from the worst response.” Additionally, explicit feedback from a judge (e.g., verifier critiques or an LLM acting as a judge) accelerates the improvement process by providing targeted guidance. Intuitively, this approach is analogous to zeroth-order optimization (Jamieson et al., 2012; Yang et al., 2024a), where two sampled values from the objective function are sufficient to guide the optimization process toward the maximum. Similarly, feedback on the best and worst responses helps steer the model toward generating responses that maximize rewards, reinforced through system prompts. (Further details in Appendix, Section B.3)
Figure 7: Text2SQL: (a)Accuracy comparison between the best-of-N method and our approach for Gemini-1.5-flash. (b) Pass@K performance comparison between queries generated using our approach and Few-shot ICL with Gemini-1.5-flash. (c) Majority@K performance comparison between queries generated using our approach and Few-shot ICL with Gemini-1.5-flash. (d) Accuracy comparison between the best-of-N method and our approach for Gemini-1.5-pro. IAD consistently outperforms all the baselines across the settings. (e) Pass@K performance comparison between queries generated using our approach and Few-shot ICL with Gemini-1.5-pro. (f) Majority@K performance comparison between queries generated using our approach and Few-shot ICL with Gemini-1.5-pro.
Key Novel Aspect of IAD : While IAD shares a high-level structure with prior approaches including Self-Refine (Madaan et al., 2023) and its variants (Chen et al., 2023; Yang et al., 2023), it fundamentally differs in the design and integration of the feedback mechanism. Prior methods like Self-Debug and ReAct use binary or execution-based signals (e.g., retry if the output fails). This is limiting when feedback is sparse, noisy or especially scenarios with sub-optimal LLM verification. These methods typically use the score as a pass/fail indicator without leveraging the full informational content of the feedback which might restricting their ability to guide clear improvements in complex tasks. In contrast, IAD is designed to extract and utilize as much signal as possible from the available feedback. Rather than relying solely on absolute reward scores, IAD transforms score-based or comparative feedback into structured guidance by identifying the best and worst responses at each iteration. These are explicitly fed back into the model through prompt conditioning, providing a clear directional signal for improvement.
Highlighting the Impact of Adaptive Feedback in IAD Beyond Sampling Gains Many prior works (Chen et al., 2023; Madaan et al., 2023) lack a clear experimental setup to disentangle the true source of improvement in their approaches. A key concern thus exists is whether the gains arise from stochastic sampling (as in Best-of-N strategies like BON), or from the actual effectiveness of self-LLM feedback. However, most prior methods do not include comparisons against stochastic sampling baselines like BON, leaving it unclear whether their feedback mechanisms meaningfully contribute to performance improvements. To address this, we conduct controlled experiment to isolate the effect of adaptive feedback in IAD from sampling diversity. Specifically, we reduce the generation temperature (to 0.1, 0.05, and 0.0), thereby minimizing randomness in the generation process. We then evaluate the performance of both IAD and BON over multiple iterations under these low-stochasticity settings as in Figures 5, 9
Key Takeaway IAD consistently improves across iterations even with minimal sampling noise, whereas BoN performance stagnates. This demonstrates that the gains from IAD are not merely due to diversity in generation, but rather from the verifier-guided adaptive feedback-driven refinement (further details in Section 4.1)
Figure 8: Performance comparison of IAD, BON with improved Model- Gemini-2.0-Flash on (a) Text2SQL and (b) Sketch2Code across varying iterations. Improvement in Multi-modal understanding of Gemini-2.0 over Gemini-1.5 as evident in performance (Baseline improves from 18 to 23). In Text2SQL, performance remains comparable to Gemini-1.5-Pro. In Text2SQL, performance of BON improves and gap reduces although still IAD outperforms.
4 Experimental Analysis
In this section, we provide detailed discussion on the experimental analysis with respect to the agentic tasks below. For all our experiments, we have utilized NVIDIA A100-SXM4-40GB GPU with 40GB of VRAM, running on CUDA 12.4 and driver version 550.90.07.
4.1 Sketch2code Results and Analysis
Evaluation metrics and Baselines. We evaluate our approach against SoTA baselines using (i) Layout Similarity, (ii) Visual IoU, and (iii) Text IoU with reference HTML (Li et al., 2024b). These metrics strongly correlate ($\approx 90\%$) with human satisfaction (Li et al., 2024b) (details in Appendix). We use layout similarity as a verifier alongside LLM-as-judge (Li et al., 2024b) to refine the generations for both BON (Beirami et al., 2024) and IAD. We compare our proposed approach IAD against single-turn SoTA models (GPT-4o, Claude-3, InternVL2, Gemini-1.5-Flash, CoT) and multi-response approaches (BON, Sk2code). We compare the performance on the same evaluation setup and dataset as used in (Li et al., 2024b).
Initial results. Figure 1 and Table 1 qualitatively and quantitatively show that single-turn approaches, including SoTA models, fail to maintain the layout structure, block positions, text placement, and sizing, leading to poor alignment with the reference layout. These models score low across all three metrics. Multi-response methods significantly improve performance, even with N=2 samples. BON, using the weaker Gemini-1.5-Flash, outperforms single-turn models. With N=4, BON surpasses SoTA single-turn models by $2-4\%$ (absolute improvement), accurately capturing block positioning, title blocks, and overall layout. Performance monotonically improves with more responses, with layout scores rising from $20.41$ to $25.7$ as N increases to six.
Benefits of IAD (ours). BON struggles with fine-grained layout details and often repeats positional errors across all N generations (Figure 1). In contrast, IAD iteratively refines responses, leading to a $\approx 3-4\%$ improvement (absolute) over BON and single-turn SoTA (Claude) in just two iterations (N=2), even with the simpler Gemini-1.5-Flash model (Table 1). At each iteration, IAD incorporates the best and worst HTML outputs as context, along with refinement instructions (with and without LLM judge). This enables progressive learning of layout components and image semantics, significantly enhancing output quality. As iterations increase, IAD achieves a high layout score of 26.75, outperforming all baselines with the same number of generations.
Method Exe Acc
DIN-SQL + GPT-4 50.72
DAIL-SQL + GPT-4 54.76
MAC-SQL + GPT-4 57.56
MCS-SQL + GPT-4 63.36
E-SQL + GPT-4o 65.58
IAD + GPT-4o 65.97
IAD + Gemini-1.5-pro 68.05
Table 2: Text2SQL - Execution accuracy comparison of previous works with our proposed approach
We also evaluate Image and Text IoU scores to check for reward over-optimization. However, as shown in Table 1 and Figure 3, text and image similarities also improve, confirming consistency in IAD’s performance gains over BON. Notably, with more generations, the performance gap between IAD and BON narrows. We also qualitatively (appendix) and quantitatively evaluate the performance across various samples with IAD which shows improvement in scores across the iterations for examples Figure 5(c). The increasing trend for samples show that our approach is able to improve the performance with increased iterations/inference compute. To further investigate and disentangle the improvement of IAD over BON or stochastic parallel sampling, we reduce the temperature (0.0, 0.05, 0.1) and perform IAD and still observe the trend in performance as shown in 5(a) 5(b), whereas the performance of BON ceases at layout score of ~21.9 even if we increase the N. This clearly shows the importance of adaptive feedback to gain a higher performance in-terms of layout score of ~26.
Figure 9: Text2SQL :Compares the accuracy of IAD on BIRD development set at low temperature across varying number of generations to highlight the improvement from iterative feedback based on IAD.
Ablation. We also consider the sensitivity of the token length of the context plays a critical role in this case, where providing the entire HTML can affect the entropy of the distribution, and thus, over-conditioning can hinder structured generation by reducing diversity and exploration. Thus, we provide only the top 200-300 tokens of the best (and worst) HTMLs. However, we remark that if there were a judge to highlight which portion of the code needs to be updated, that would be more targeted. Hence, we incorporate LLM-judge (Gemini-1.5-Pro), which has the reference policy, and it checks with the current response and provides feedback on improvement and sometimes snippets of HTML as well (however, we restrict that to 200 tokens -5-6% of the original HTML). This leads to an additional improvement of  2-4% (absolute improvement) for the layout score with just two iterations and a final score of ~28 with 6 iterations, demonstrating the importance of iterative approaches with judge feedback for agent performance.
Figure 10: Qualitative illustration of IAD in improving the action taking capabilities of Webagent for Webshop environment (Yao et al., 2023) over direct generation, where IAD improves the search action by iteratively reviewing and refining the actions.
We have kept the judge-prompt similar as Sk2code (Li et al., 2024b) for fairness in evaluation. It is important to note that IAD-fb represents that we augment the judge feedback along with our best and worst response and generic instruction prompt as described in Algorithm. We admit that its indeed possible to learn more specific prompts to construct a better judge via training/prompt-tuning, however that’s outside the main scope of this work. Overall, in all our analyses, our findings remain consistent, where IAD outperforms baselines.
4.1.1 Text-to-SQL Results and Analysis
In this section, we detail the experiments conducted on the BIRD text-to-SQL benchmark (Li et al., 2024a). Our evaluation involved two models, Gemini-1.5-pro and Gemini-1.5-flash, to test our proposed approach and compare it against baseline methods.
IAD vs test-time-compute methods:
We compared our method to two baselines: few-shot chain-of-thought prompting and the best-of-N approach, both implemented using the same models. These comparisons highlight the effectiveness of our multi-turn strategy in narrowing the gap between the reference distribution and the optimal distribution. Our multi-turn approach addresses common issues in single-turn generation, which can produce SQL queries with syntax and semantic errors. By incorporating multiple rounds of response selection and feedback, our method effectively corrects these errors. Moreover in the multi-turn setting, the reward model benefits from observing query execution results, allowing it to more readily identify syntax and semantic issues. For the few-shot prompting baseline, we employed two metrics—Pass@K and Majority@K—to evaluate the quality of the generated candidate queries. In order to obtain candidate queries we execute our approach multiple times with three rounds of refinement. As shown in Figure 7, our approach consistently produces candidate sets with higher correctness. In contrast, the best-of-N method aggregates results, using self-consistency; therefore, we used the original execution accuracy for its evaluation. To ensure a fair comparison, we computed accuracy across different numbers of LLM calls. In our method, we first generate a draft SQL query, followed by multiple rounds of evaluation, feedback generation, and query refinement, each contributing to the LLM call count. As illustrated in Figure 7, our approach consistently produce higher accuracy across different number of LLM calls compared to the best of N approach. Similar to the previous section, Figure 9 demonstrates the performance improvement when increasing the number of LLM calls across different temperatures.
IAD vs Text-to-SQL baselines
We also compare our approach with several baseline methods in the text-to-SQL domain that do not rely on fine-tuning on the BIRD train set, ensuring a fair comparison. These baselines originate from different families of methods, for example, MCS-SQL (Lee et al., 2024) is a variant of the best-of-N method, DIN-SQL (Pourreza and Rafiei, 2024) and DAIL-SQL (Gao et al., 2023) employ human-engineered few-shot prompts to achieve higher performance, and MAC-SQL (Wang et al., 2023) is a multi-agent approach designed specifically for text-to-SQL tasks. The results of these comparisons are provided in Table 2, where we achieved the highest execution accuracy by 3% margin compared to the second best method.
4.2 Webshop Results and Analysis
Models (PR) (SR)
Lemur-70b 71 11
Mistral-7b 68.2 13.9
Vicuna-13b-16k 73 21
Gemini-1.5-Flash 71.3 26.5
GPT-3.5-Turbo-16k 73 27
Text-Davinci-003 72 29
Gemini-1.5-Pro 73.5 29.3
BON-SC + Gemini-1.5-Pro 74.12 30.31
DeepSeek-67b 72 31
GPT-3.5-Turbo 76 35
IAD + Gemini-1.5-Pro 71 38.3
GPT-4 75.8 38.5
GPT-4o 73.1 40.3
BON-SC + GPT-4o 74.21 41.09
IAD + GPT-4o 74.6 44.68
Table 3: Webshop- Progress Rate (PR) and Success Rate (SR) for Models in the Webshop Environment (Yao et al., 2023). Perform a comparison of IAD (with two models GPT-4o and Gemini-1.5-Pro) against SoTA baselines for the Webshop Leaderboard with the evaluation similar to followed in Agentboard (Ma et al., 2024). For BON-SC, we generated N=3 responses, similarly we used three feedback steps for IAD.
For all the experiments, we use greedy decoding with temperature=0.1, top p = 1.0 for our results. For BON-SC, we try to vary the temperature to generate more diverse responses. We compare IAD* against SoTA baselines, including Lemur, Mistral, Vicuna, Gemini, and GPT models, in the Webshop environment, where agents interact with webpages through search and click actions. Performance is measured via progress rate (PR) and success rate (SR), with SR being the key metric for exact attribute satisfaction. We follow the same evaluation methodology as AgentBoard (Ma et al., 2024). IAD* outperforms baselines, achieving  8-10% improvement (absolute) in SR across most models with Gemini-1.5 inclduing BON and achieves an improvement of over 3-4% (absolute) with very strong GPT-4o as shown in Table 3.
To further enhance performance, we introduce feedback-based iterative refinement, where agents receive prompts like "Reflect on your previous action and improve upon it", "Improve the previous actions focusing on key-attributes", "Follow the format to be action.name-’search’ and action.params -’red tshirt….’" for allowing LLMs to iteratively refine their decisions. A verifier-based LLM judge determines when to iterate or finalize actions. We also compare IAD* with BON-SC (Best-of-N with Self-Consistency, N=3) but find that BON provides limited improvement due to low diversity in generations, even with temperature = 0.6. Beyond this, higher temperatures introduce irrelevant attributes, reducing effectiveness.
Figure 11: Webshop : Progress-rate comparison of IAD over Best of N with Self-consistency (BON-SC) and baseline on random selected 20% of the total evaluation set using weaker Gemini-1.5-Flash-01 (weaker model). Figure highlights the importance of iterative feedback while generation. Evaluation was done with a weaker model and on a subset of eval examples.
In contrast, IAD* improves search queries by refining unwanted attributes, enhancing focus on key parameters. Additionally, to show the effectiveness, we perform with a weaker version of Gemini-1.5-Flash-001 and perform the experiment, where we observe that it needs more iterations to actually follow the action format, without this iteration it fails to provide response in a meaningful format. On the other-hand, we approach the same with BON-SC which couldn’t improve much due to the issue in diverse response generation as highlighted before.
5 Conclusion
In this work, we proposed IAD : an iterative decoding approach for AI agent alignment with black box access which highlights the effectiveness of iterative decoding (guided by a verifier) for these complex agentic tasks. IAD enhances prior baselines by dynamically integrating verification and critique mechanisms to iteratively refine and improve response quality. Experimental results across diverse domains show that our approach consistently outperforms several existing baselines including highly competitive BON highlighting its ability to bridge the gap between reference and optimal policies at inference time.
6 Limitations
While IAD — our proposed iterative decoding approach — improves upon prior baselines by leveraging verifier feedback, it is inherently sequential, leading to increased computational overhead compared to easily parallizable BON approaches. Addressing this trade-off between quality improvement and computational efficiency remains an important area for future research. Efficient optimization approaches, such as adaptive stopping, speculative decoding, and more efficient verifier-guided selection, could improve the efficiency in iterative decoding for agentic tasks. Additionally, verifier and judge plays a crucial role in our approach. Thus a more concrete investigation and selection of a judge for these challenging tasks is a valid and crucial next step of our work. We highlight that this work is of academic nature and has no direct or immediate harmful impacts to society. However, since this work deals with improving AI agents, it should be done under safety protocols and guidelines. We want to highlight that this study is limited to English language text primarily due to the nature of open-source datasets used.
[…]
Appendix A Detailed Environment Description
1. Text-to-SQL Text-to-SQL serves as a critical interface between natural language and structured query languages by enabling users to translate natural language queries into executable SQL commands. This functionality empowers individuals without SQL expertise to interact with complex databases, thereby facilitating data exploration, informed decision-making, automated analytics, and advanced feature extraction for machine learning. Generally, a Text-to-SQL system receives a natural language question and any pertinent metadata about the tables and columns, which serves as external knowledge to aid in database comprehension. Consequently, such systems are responsible not only for interpreting user intent and identifying relevant information from a potentially vast set of tables and columns but also for generating SQL queries that may include multiple conditions—a process that is inherently reasoning intensive. To evaluate our proposed framework, we employ the BIRD benchmark (Li et al., 2024a), a challenging and widely used dataset in the Text-to-SQL domain. BIRD comprises an extensive collection of 12,751 unique question-SQL pairs drawn from 95 large databases with a total size of 33.4 GB. The benchmark spans more than 37 professional domains, including blockchain, hockey, healthcare, and education, making it a comprehensive resource for assessing the robustness and generalizability of Text-to-SQL systems. The primary metric for model comparison in this domain is execution accuracy (EX), where the ground truth SQL query and the predicted SQL query are both executed over the target database, if they both generate same sets of results the accuracy for the predicted SQL query is considered as accurate.
(a) Sketch2Code Environment
(b) Text-to-SQL Environment
(c) Webshop Environment
Figure 12: These three figures given an overview of three diverse and challenging agentic tasks that we consider to evaluate the performance of agents with our proposed approach vs baselines -(a) Sketch2code(Li et al., 2024b) (b)Text2SQL (Li et al., 2024a) and (c) Webshop (Yao et al., 2023)
2. Sketch2code: Sketch2code (Li et al., 2024b) challenges and evaluates the multi-modal capabilities of agent where the objective is transform wireframe-style rough userk sketches into functional HTML prototypes with embedded CSS. Sketch2Code uniquely tests multi-modality, requiring structured code generation from imprecise visual input, often leading to misaligned text, incorrect spacing, and structural inconsistencies. This leads to challenges such as misaligned text, incorrect spacing, missing components, structural inconsistencies, making it an extremely challenging benchmark for multimodal LLMs. The complexity of this task arises from: ambiguity in hand-drawn sketches, where component boundaries, spacing, and positioning are not precisely defined. The evaluation of the generation is done primarily with three key metrics : Layout Similarity, Text IOU, Image IOU. Layout Similarity (IoU-based metrics): Intersection-over-Union (IoU) is computed for different UI components (e.g., buttons, images, text blocks) to measure how well their positions match the reference. Intersection-over-Union (IoU) is computed for different UI components (e.g., buttons, images, text blocks) to measure how well their positions match the reference implementation. Text-IOU similarly measures how accurately the generated text aligns with the reference design. Image IOU uses CLIP embeddings to compare the visual appearance of the generated webpage with the reference design and evaluates color similarity, element positioning, and component rendering. These metrics provide a reliable way to measure the quality of the generated response and strongly correlates with human judgement. Evaluations are done also with LLM as judge to compare the performance.
3. Webshop is a large-scale, web-based interactive environment designed to test an AI agent’s capability to perform sequential decision-making in an online shopping scenario under sparse feedback (Yao et al., 2023). The environment is modeled as a partially observable Markov decision process, where the agent navigates a simulated e-commerce platform to fulfill a user’s product request based on natural language instructions. At each step, the agent receives an observation in the form of a webpage—such as search results, product details, or checkout options—and must decide on an action, including searching for a product, clicking on an item, or selecting options. The evaluation is based on success rate (SR), which measures whether the agent successfully selects a product that matches all specified criteria (attributes, price, and options), and task score, which represents the overall alignment of the final selection with the given instruction. The WebShop environment presents significant challenges, including sparse rewards (since feedback is only provided at the end of an episode), the need for strategic backtracking and exploration, and handling noisy or ambiguous natural language instructions. This setup makes WebShop a rigorous benchmark for evaluating long-horizon reasoning, language understanding, and decision-making in real-world-like online navigation scenarios.
A.1 Limitation of Single-turn Approach
In this section we characterize the performance gap $\Delta$ as the difference between the reward the optimal or ground-truth agent is achieving vs the reward achieved by the reference achieved by the reference agent policy (Mroueh, 2024):
$\Delta$ $=\mathbb{E}_{y\sim\pi^{*}(\cdot|x)}[R(x,y)]-\mathbb{E}_{y\sim\pi_{0}(\cdot|x)}[R(x,y)]$
$\leq\sup_{R\in\mathcal{R}}\mathbb{E}_{y\sim\pi^{*}(\cdot|x)}[R(x,y)]-\mathbb{E}_{y\sim\pi_{0}(\cdot|x)}[R(x,y)]$
$\leq\|R\|_{\text{max}}d_{\text{TV}}(\pi^{*}(\cdot|x),\pi_{0}(\cdot|x)),$
where $R(x,y)$ represents the reward function measuring the quality of the generated response, and $d_{\text{TV}}(\pi^{*}(\cdot|x),\pi_{0}(\cdot|x))$ is the total variation (TV) distance between the optimal policy $\pi^{*}(\cdot|x)$ and the reference policy $\pi_{0}(\cdot|x)$. This result demonstrates that the performance gap $\Delta$ is inherently limited by the quality of the reference agent policy $\pi_{0}(\cdot|x)$, as measured by its divergence from the optimal policy. Thus, if $\pi_{0}(\cdot|x)$ is close to $\pi^{*}(\cdot|x)$ (in terms of TV distance), the performance gap will be small, resulting in near-optimal responses and viceversa.
Figure 13: Sketch2code: Qualitative evaluation of the generated HTMLs with BON sampling (N=4) corresponding to the user-sketch (left-bottom) and reference html (left-top). The figures show that BON performs much better in matching the reference HTML but still misses specific properties like rectangular structue, position of text, relative positioning of smaller blocks etc.
Appendix B Detailed Experimental Analysis
B.1 Text-to-SQL Detailed Results and Analysis
In this section, we detail the experiments conducted on the BIRD text-to-SQL benchmark (Li et al., 2024a). For these experiments, we employed the Gemini-1.5-pro and Gemini-1.5-flash models both to generate actions at each state and as judge models to predict the reward. At each state, the LLM is provided with the database schema and the user’s query, based on which it generates a draft SQL query. This draft query is then evaluated by the judge model, which also produces feedback on how to improve the draft. The LLM uses this feedback to generate a revised query, establishing a self-correction loop. Finally, the answer with the highest reward value is selected as the candidate output. This process can be repeated to generate multiple candidate SQL queries. We then apply self-consistency (Wang et al., 2022) by executing all candidate queries over the database, grouping them based on their execution results, and selecting a query from the largest result cluster as the final answer. We first compare our proposed method with the widely used few-shot prompting approach in terms of Pass@k performance and final accuracy after self-consistency (Majority@K) using execution accuracy as the metric in order to demonstrate that using our method we can generate a pool of candidates with a higher quality. Subsequently, we compare our approach with the best-of-N approach which is one of the strong baselines as a test-time compute approaches to demonstrate the effectiveness of the proposed framework. Finally, we compare our method with all previously proposed test-time methods on the BIRD development set benchmark, excluding works that rely heavily on fine-tuning LLMs (Pourreza et al., 2024; Talaei et al., 2024; Maamari et al., 2024; Gao et al., 2024) for a fair comparison.
Figure 14: Sketch2code: Provides a qualitative verification of layout score as a metric and corresponding correlation to human judgement. It is evident that HTMLs with higher match with the reference layout (right-top) and user sketch(right-bottom) has higher layout score and vice-versa showing that its a valid metric.
[t] Sketch2code : Oracle Judge Prompt for providing Feedback Judge Prompt for Sketch2code (similar to (Li et al., 2024b)): Act as you are a front-end designer working with a code agent to implement an HTML webpage . You are provided with two HTMLs : the first is the reference webpage, and the second one is the current implementation from the code agent . The task is to carefully compare the agent ’s implementation against the reference webpage , and provide feedback to help the agent make its implementation closer to the reference webpage . Your feedback should be specific to the differences in layouts and visual components on the two webpages. Don’t focus on the style components too much and focus on layout similarity and visual match with the reference webpage.
[t] Sketch2code : Feedback from LLM as a Judge (Self Verification) Feedback provided from Self-LLM : Iter1 : The HTML structure and CSS styling do not reflect the provided wireframe. Iter2 :The layout needs to be revised to accurately represent the sketch’s two-column image section and the distribution of text blocks. Iter 3: The layout of the text blocks and image containers does not accurately reflect the provided wireframe. The layout uses flexbox but doesn’t accurately reflect the sketch’s proportions and image placement. The large image should be centered and the smaller images should be positioned to the left of their respective text blocks. Iter4 : Implement a more precise grid-based layout using CSS grid or flexbox to achieve the correct positioning and sizing of all elements. Iter 5: The provided HTML closely resembles the wireframe but still needs significant layout adjustments. Use CSS Grid to precisely position and size all elements according to the wireframe’s proportions.
B.2 Sketch2code
For Sketch2code (Li et al., 2024b), we provide a detailed comparison of our approach against SoTA baselines on several evaluation criterion and metrics. We used the hyperparameter setting of temperature = 0.6, max tokens = 4096, top p = 1.0, frequency/repetition penalty = 0.0, and presence penalty = 0.0 for all our results. For the metrics, we consider metrics centring 1.Layout Similarity, 2. Visual IoU, 3. Text IoU with reference HTML following (Li et al., 2024b). These metrics offer a comprehensive and reliable assessment of HTML generation quality, demonstrating a strong correlation (9̃0%) with human satisfaction, as shown in (Li et al., 2024b) (further details in Appendix). Hence, we use Layout similarity as a verifier along with LLM-as-judge (Li et al., 2024b) to guide the generations for both BON (Beirami et al., 2024) and IAD. We report comparison with baseline single-turn approaches including SotA models GPT-4o, Claude-3, InternVL2, Gemini-1.5-Flash, CoT and variants along with multi-response generation approaches including BON, Sk2code and IAD (Ours). Single turn approaches even from SoTA models fail to match the layout structure, position of blocks, textual content, size of the blocks etc in the given user-sketch, causing a mismatch w.r.t to the reference layout as can be clearly seen in Figure and achieves a low score in-terms of all the three metrics in-comparison with multi-response generation approaches even with N=2. Best-of-N sampling (BON) with a weaker model Gemini-1.5-Flash improves over single-turn approaches and , with $N=4$ generations, it outperforms SoTA models with single-turn responses by a margin of 15-18%, by correctly identifying the block position, title block, overall layout structures etc. We see monotonic improvement in performance over the number of responses as the layout score improves from $20.41$ to $25.7$ with 6 responses. However, BON struggles in incorporating fine-grained details about layout structure and makes some-times makes repetitive mistakes in the position of block in all the $N$ generation for the prompts (as shown in Fig). Our proposed approach IAD, mitigates this gap by iteratively improving the responses and as shown in Table 1, it achieves a major improvement of $\leavevmode\nobreak\ 3-5\%$ (absolute improvement) from BON as well as single-turn SoTA Claude with just 2 iterations (eq : N=2) even with simpler model Gemini-1.5-Flash. At each iteration, we pass the best and worst HTML as a context along with instruction, for generating the next iteration. We observe IAD is able to learn fine-grained layout components, image semantics over iterations with the context of the Best and Worst HTML. We see that with increased iterations, performance of IAD improves reaching to a very high layout score of 26.75, outperforming all baselines with same generations. We also report the Image and Text IoU scores while optimizing with the layout-score, to check for reward-overoptimization of the metric.
Figure 15: Sketch2code : Top row shows the user sketch, reference image and the performance of IAD over iterations. The figure highlights improvement of IAD over 4-turns w.r.t Layout similarity score (1/100) for 3 examples. It shows clear improvement over iterations. We also qualitatively analyse the snapshots of the HTMLs generated by the agent, which demonstrates that over iteration the qualitative performance improves and matches the input sketch/reference HTML.
However, as can be observed in Table 1 and Figure-3, that text and image similarities are also improving over the iterations and our findings regarding comparison with baseline BON are consistent with the same. However, we observe that with increase number of generations the performance gets closer to BON. We also consider sensitivity of the token-length of the context plays a critical role in this case, where providing the entire HTMLs can affect the entropy of the distribution, and thus over-conditioning can hinder structured generation by reducing diversity and exploration (as shown in Figure ). Thus, we provide only the top 200-300 tokens of the best (and worst) HTMLs. However, it is clear that if there would be a judge to highlight which portion of the code needs to be updated that will be more targeted. Hence, we incorporate LLM-judge (Gemini-1.5-Pro) which has the reference policy and it checks with the current response and provide feedback on improvement and sometimes snippets of HTML as well (however, we restrict that to 100 tokens   5-8% of the original HTML). This leads to an additional improvement of  2-4% for the layout score with just two iteration and final score of  28 with 6 iterations, demonstration the important of iterative approaches for agent performance. We have kept the judge-prompt similar as Sk2code (Li et al., 2024b) for fairness in evaluation. We admit that its indeed possible to learn more specific prompts to construct a better judge via training/prompt-tuning, however thats outside the main scope of this work. Overall, in all our analyses, our findings remain consistent, where IAD outperforms baselines.
Figure 16: Sketch2code: (a)Represents the performance of IAD w.r.t Layout score with varied sparsity NS (No Sparsity), LS (Low Sparsity) MS (Med Sparsity) and HS (High Sparsity) (b) Performance comparison with BON via varying sparsity which shows the improvement over BON is less with sparsity (c) Performance of IAD w.r.t Layout similarity with Sparsity via across varying the number of generations.
B.3 Importance of Verifier and Reward functions
In this section, we provide a motivation and importance of the verifier in ensuring monotonic improvement with our proposed approach. We define $\pi^{*}(\cdot|x)$ as the target policy generating $y^{*}\sim\pi^{*}(\cdot|x)$. At each step $t$, we sample $y_{t+1}\sim\pi_{0}(\cdot|x,\hat{y}_{t})$, where $\hat{y}_{t}$ is the best response so far, and update $\hat{y}_{t+1}=\arg\max{y\in(y_{t},\hat{y}_{t})}R(x,y)$, accepting $y{t+1}$ if $R(x,y_{t})-R(x,\hat{y}_{t})>0$. If $R(x,y)$ incorporates the information of $\pi^{*}(y|x)$ (upto a normalization i.e $R(x,y)=\frac{f(\pi^{*}(y|x))}{Z}$, $f$ being a monotonic function), we show that our iterative refinement never deteriorates performance. In other words, we assume that for any two responses $y_{1},y_{2}$, if the reward function (Setlur et al., 2025) satisfies $R(x,y_{1})>R(x,y_{2})$ then it implies that the optimal policy assigns a higher probability to $y_{1}$ than $y_{2}$, i.e $\pi^{*}(y_{1}|x)>\pi^{*}(y_{2}|x)$. A natural way to measure closeness to the optimal response is by estimating the distance under the true probability distribution (i.e target policy) $\pi^{*}(\cdot|x)$, defined as
$d(\hat{y}_{t+1},y^{*})=\pi^{*}(y^{*}|x)-\pi^{*}(\hat{y}_{t+1}|x)$ (4)
where the difference captures that how good the quality of the response is under optimal policy. If the response $\hat{y}_{t+1}$ is highly optimal, then $d(\hat{y}_{t+1},y^{*})$ will be low and viceversa, when $\hat{y}_{t+1}=y^{*}$, the gap will be zero.
$d(\hat{y}_{t},y^{*})$ $=\pi^{*}(y^{*}|x)-\pi^{*}(\hat{y}_{t}|x)$ (5)
$=\pi^{*}(y^{*}|x)-\pi^{*}(\hat{y}_{t}|x)-(\pi^{*}(\hat{y}_{t+1}|x)-\pi^{*}(\hat{y}_{t}|x))$
$\leq\pi^{*}(y^{*}|x)-\pi^{*}(\hat{y}_{t}|x)=d(\hat{y}_{t},y^{*})$
where, we first add and subtract the term $\pi^{*}(\hat{y}_{t}|x)$. Then by definition of our acceptance rule, we ensure that $\pi^{*}(\hat{y}_{t+1}|x)-\pi^{*}(\hat{y}_{t}|x)\geq 0$, where equality occurs when $\hat{y}_{t+1}=\hat{y}_{t}$. Thus we have $d(\hat{y}_{t})<=d(\hat{y}_{t-1})$ i.e we ensure that the responses over the iteration are either improving or remain the same over iteration and won’t deteriorate over iterations. However, it is important to note that this is based on the assumption that the reward function is aligned with the optimal distribution, meaning that selecting responses based on maximizing $R(x,y)$ leads to responses that are increasingly closer to the ground-truth distribution $\pi^{*}(\cdot|x)$. This highlights the criticality and accuracy of a true verifier in these agentic tasks as also shown in (Setlur et al., 2025)
Verifier and Reward function: We provide qualitative evaluation of considering layout similarity as a verifier due to its Interpretability and also correlation with human judgements also shown in (Li et al., 2024b). Additionally, we want to highlight that Sketch2code represents an extremely complex and challenging task for using self-LLM as a judge (Madaan et al., 2023) (without significant prompting) to compare between two generated HTMLs (by the agent) with its similarity to the input sketch and prompt. The input sketch has entirely different distribution than the image snapshot of the generated HTML which makes it harder for LLM as a judge to perform which is one of the reason we hypothesize that Self-refine (Madaan et al., 2023) type approaches doesn’t provide improvements as shown in Table 1. On the other-hand, although LLM judge (oracle) provides more meaningful feedback when it has access to the reference HTML, however needs to be prompted efficiently to generate meaningful responses. We agree the fact that our judge (oracle) for the feedback was allowed to provide more context than the one used in (Li et al., 2024b). However, the performance improvement in (Li et al., 2024b) feedback approch is very less and we hypothesize major reasons can be not performing IAD type approach, where we take previous best response (HTML) in the context along with specific instructions. Even for LLM-judge (oracle), we leverage feedback along with the previous best and worst HTMLs, which helps in providing more meaningful context to the agent in generating the correct HTML.
Figure 17: Text2SQL: An example of two responses is presented: the first response, generated using our proposed approach, is correct, while the second response, produced using the best-of-N method, is incorrect.
Query Search Attempts Results Found Final Outcome
Men’s Black Loafers (Size 10.5, Rubber Soles, <60) Multiple searches, clicked "Next" repeatedly, found unrelated shoes (sneakers, sandals, pumps) None matched the requirement Task Failed - No suitable options found (Reward: 0.0)
Blue Diamond Almonds (Gluten-Free, Pecan, 12 Pack) Repeated searches, encountered "No Search button" error multiple times, retrieved irrelevant snack items Nut Thins Crackers, Keto Bars, M&M’s Chocolate Task Failed - No relevant product found (Reward: 0.0)
Folding Storage Box Ottoman (Faux Leather, 60x40x40cm, <170) Initial product matched but had incorrect size, next searches returned irrelevant furniture items Found an ottoman, but wrong size & overpriced Task Failed - No exact match found (Reward: 0.0)
Official Cleveland University Drawstring Shorts (Small, Charcoal, Machine Washable, <60) Search led to incorrect results (Marvel T-Shirts, Women’s Yoga Shorts), agent attempted refinement but couldn’t find exact product No official Cleveland University shorts found Task Failed - No suitable options found (Reward: 0.0)
Organic Hair Growth Serum Roller Set (For All Hair Types, <60) Search retrieved some serums but none matched exact request (wrong quantity or expensive) Found a set, but incorrect product version Task Failed - No exact match found (Reward: 0.0)
Table 4: Webshop : Highlights several Failure Cases of the Baseline Agent (Gemini-1.5-Pro) in Retrieving Relevant products given the task. This represents the challenge of current model in performing strategic exploration in Webshop.
[…]
Webshop - Task: Buy a Folding Storage Box Ottoman- IAD (Success) Size: 60x40x40cm  Material: Faux Leather  Price: Under \$170 Search $\rightarrow$ "folding storage box ottoman faux leather 60x40x40cm" Product List $\rightarrow$ Found 50 results Ottoman Footstool (40x40x40cm) - \$149.97 Other options did not match size or price Click $\rightarrow$ Select "Ottoman Footstool" Size Selection $\rightarrow$ Click "60x40x40cm" Buy Now $\rightarrow$ Proceed to checkout Task Completed
Webshop - Task: Buy a Vegan, Gluten-Free Protein Shake - IAD (Success) Requirements: 100% Vegan, Gluten-Free, Soy-Free  Price: Under \$40 Search $\rightarrow$ "gluten free vegan plant based protein shake" Product List $\rightarrow$ Found 50 results OWYN Protein Shake (Cold Brew Coffee, 12oz) - \$11.07 Other products exceeded price or dietary restrictions Click $\rightarrow$ Select "OWYN Protein Shake" Buy Now $\rightarrow$ Proceed to checkout Task Completed
B.4 Text2SQL
In this section we provide some example responses of Best of N approach and our proposed method on the BIRD development set samples.
[t] Text2SQL : Generated queries for best of N approach (Wrong answer) DB schema: CREATE TABLE satscores ( enroll12 INTEGER not null, dname TEXT null, ); CREATE TABLE frpm( IRC INTEGER null, ‘Free Meal Count (K-12)‘ REAL null, ); CREATE TABLE schools( GSserved TEXT null, Street TEXT null, ); Question: Which active district has the highest average score in Reading? Predicted Query: SELECT T1.District FROM schools AS T1 INNER JOIN satscores AS T2 ON T1.CDSCode = T2.cds WHERE T1.StatusType = ’Active’ GROUP BY T1.District ORDER BY AVG(T2.AvgScrRead) DESC LIMIT 1
[t] Text2SQL : Generated queries for our method (Correct answer) DB schema: CREATE TABLE satscores ( enroll12 INTEGER not null, dname TEXT null, ); CREATE TABLE frpm( IRC INTEGER null, ‘Free Meal Count (K-12)‘ REAL null, ); CREATE TABLE schools( GSserved TEXT null, Street TEXT null, ); Question: Which active district has the highest average score in Reading? Predicted Query: SELECT T1.dname FROM satscores AS T1 INNER JOIN schools AS T2 ON T1.cds = T2.CDSCode WHERE T2.StatusType = ’Active’ AND T1.AvgScrRead IS NOT NULL GROUP BY T1.dname ORDER BY MAX(T1.AvgScrRead) DESC LIMIT 1
[t] Text2SQL : Generated queries for best of N approach (Wrong answer) DB schema: CREATE TABLE satscores ( enroll12 INTEGER not null, dname TEXT null, ); CREATE TABLE frpm( IRC INTEGER null, ‘Free Meal Count (K-12)‘ REAL null, ); CREATE TABLE schools( GSserved TEXT null, Street TEXT null, ); Question: Among the schools with the average score in Math over 560 in the SAT test, how many schools are directly charter-funded? Predicted Query: SELECT COUNT(T1.CDSCode) FROM schools AS T1 INNER JOIN satscores AS T2 ON T1.CDSCode = T2.CDSCode WHERE T2.AvgScrMath > 560 AND T1.FundingType = ’Directly funded’
[t] Text2SQL : Generated queries for our method (Correct answer) DB schema: CREATE TABLE satscores ( enroll12 INTEGER not null, dname TEXT null, ); CREATE TABLE frpm( IRC INTEGER null, ‘Free Meal Count (K-12)‘ REAL null, ); CREATE TABLE schools( GSserved TEXT null, Street TEXT null, ); Question:Among the schools with the average score in Math over 560 in the SAT test, how many schools are directly charter-funded? Predicted Query: SELECT COUNT(DISTINCT T1.CDSCode) FROM schools AS T1 INNER JOIN satscores AS T2 ON T1.CDSCode = T2.cds INNER JOIN frpm AS T3 ON T3.CDSCode = T2.cds WHERE T2.AvgScrMath > 560 AND T3.‘Charter Funding Type‘ = ’Directly funded’ AND T3.‘Charter Funding Type‘ IS NOT NULL

### v4

@@
-Review, Refine, Repeat: Understanding Iterative Decoding of AI Agents with Dynamic Evaluation and Selection
+On the Role of Feedback in Test-Time Scaling of Agentic AI Workflows
 Abstract
-While AI agents have shown remarkable performance at various tasks, they still struggle with complex multi-modal applications, structured generation and strategic planning.
-Improvements via standard fine-tuning is often impractical, as solving agentic tasks usually relies on black box API access without control over model parameters.
-Inference-time methods such as Best-of-N (BON) sampling offer a simple yet effective alternative to improve performance.
-However, BON lacks iterative feedback integration mechanism.
-Hence, we propose Iterative Agent Decoding (IAD) which combines iterative refinement with dynamic candidate evaluation and selection guided by a verifier.
-IAD differs in how feedback is designed and integrated, specifically optimized to extract maximal signal from reward scores.
-We conduct a detailed comparison of baselines across key metrics on Sketch2Code, Text2SQL, and Webshop where IAD consistently outperforms baselines, achieving 3–6% absolute gains on Sketch2Code and Text2SQL (with and without LLM judges) and 8–10% gains on Webshop across multiple metrics.
-To better understand the source of IAD’s gains, we perform controlled experiments to disentangle the effect of adaptive feedback from stochastic sampling, and find that IAD’s improvements are primarily driven by verifier-guided refinement, not merely sampling diversity.
-We also show that both IAD and BON exhibit inference-time scaling with increased compute when guided by an optimal verifier.
-Our analysis highlights the critical role of verifier quality in effective inference-time optimization and examines the impact of noisy and sparse rewards on scaling behavior.
-Together, these findings offer key insights into the trade-offs and principles of effective inference-time optimization.
+Agentic AI workflows (systems that autonomously plan and act) are becoming widespread, yet their task-success rate on complex tasks remains low.
+A promising solution is inference-time alignment, which uses extra compute at test time to improve performance.
+Inference-time alignment relies on three components: sampling, evaluation, and feedback.
+While most prior work studies sampling and automatic evaluation, feedback remains under-explored.
+To study the role of feedback, we introduce Iterative Agent Decoding (IAD), a procedure that repeatedly inserts feedback extracted from different forms of critiques (reward models or AI-generated textual feedback) between decoding steps.
+Through IAD, we analyze feedback along four dimensions: (1) its role in the accuracy–compute trade-offs with limited inference budget, (2) quantifying the gains over diversity-only baselines such as best-of-N sampling, (3) effectiveness of composing feedback from reward models versus textual critique, and (4) robustness to noisy or low-quality feedback.
+Across Sketch2Code, Text2SQL, Intercode, and WebShop, we show that IAD with proper integration of high fidelity feedback leads to consistent gains up to 10% absolute performance improvement over various baselines such as best-of-N.
+Our findings underscore feedback as a crucial knob for inference-time alignment of agentic AI workflows with limited inference budget.
 1 Introduction
-AI agents have demonstrated remarkable capabilities across a wide range of tasks, including natural language understanding, code generation, and mathematical reasoning (Fu et al., 2024; Liu et al., 2024; Jiang et al., 2024; Patel et al., 2024; Huang et al., 2024; Liu et al., 2023).
-However, state-of-the-art AI agents still face challenges when handling tasks that require understanding multimodal inputs and reasoning through structured output generation (Costarelli et al., 2024; Kinniment et al., 2024; Liu et al., 2023; Verma et al., 2024; Valmeekam et al., 2024).
-For example, consider agentic tasks such as Sketch2Code (Li et al., 2024b), Text2SQL (Li et al., 2024a), and the web-navigation task Webshop (Yao et al., 2023).
-Sketch2Code requires an AI agent to generate structured HTML from rough sketches, often resulting in incorrect generations.
-Text2SQL involves translating natural language queries into executable SQL statements, demanding precise reasoning while avoiding semantic and logical errors.
-Meanwhile, Webshop is a sequential decision-making task that evaluates an agent’s ability to navigate an e-commerce platform effectively.
-Across these tasks, AI agents consistently underperform compared to human experts, underscoring the limitations of current agentic models in handling complex, structured outputs and sequential reasoning (Yao et al., 2023; Liu et al., 2024; Li et al., 2024b) where agents achieve merely 20-30% performance in task specific metrics.
-Figure 1: Qualitative (Left) and Quantitative (Right) illustration of the performance benefit of IAD (Ours) over Single-turn response generations using Gemini-1.5 (Base) Text2SQL task (Li et al., 2024a).
-IAD improves performance by correctly handling query logic and joins, improving the accuracy over baseline and BON (Best-of-N).
-A key challenge in improving agentic AI systems is their ability to handle structured reasoning and multimodal tasks effectively.
-One promising approach to addressing this challenge is fine-tuning a new model through alignment (Ouyang et al., 2022a).
-A key alignment technique, reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022a, b; Bai et al., 2022) has significantly enhanced the capabilities of generative models, it is not directly applicable to AI agents due to their inherently black box nature.
-Most agentic systems and frameworks (Liu et al., 2024; Fu et al., 2024) operate in environments where internal model updates are infeasible, making fine-tuning-based alignment techniques inapplicable.
-An alternative approach is inference-time alignment, such as controlled decoding (Khanov et al., 2024; Deng and Raffel, 2023; Mudgal et al., 2024; Han et al., 2024; Chakraborty et al., 2024; Yang and Klein, 2021), which can align generative models without fine-tuning.
-However, these methods still rely on access to token-level logits from the reference model to steer responses, an assumption that does not hold for black box agentic systems.
-These limitations highlight the need for an inference-time alignment technique that is (i) compatible with various underlying models (without requiring model-specific adaptations), (ii) easy to implement (requires no modification of the underlying models), and (iii) commercial API-friendly (Applicable to closed-source models that are only accessible via API.
-This leads us to an important question:
-How can we improve the inference alignment of AI Agents with black box access for complex and structured tasks?
-A simple solution to the above question is Best-of-N (BON) sampling and its variants (Nakano et al., 2021; Stiennon et al., 2020; Beirami et al., 2024), as it is black box compatible and logit-free, making it a natural choice for inference-time alignment.
-BON selects the best response from multiple generated candidates using a verifier, improving response quality in structured tasks such as HTML generation and SQL query synthesis.
-BON is even almost optimal in terms of quality (Gui et al., 2025; Yang et al., 2024b; Beirami et al., 2024; Amini et al., 2024).
-However, BON may be ineffcient requiring a large N to achieve reasonable performance.
-To address these limitations, we propose Iterative Agent Decoding (IAD)—an iterative approach that refines responses through verifier-guided feedback.
-IAD differs fundamentally in the design and integration of its feedback mechanism, specifically optimized to extract and utilize rich signals from the reward score thus demonstrating the effectiveness of adaptive feedback in complex tasks.
-We then conduct a detailed comparison of baselines including parallel sampling (BON) and iterative decoding (IAD) across several key metrics relevant to inference-time scaling, yielding key insights into the trade-offs and principles of effective inference-time optimization.
-We summarize our contributions and findings as follows.
-Figure 2: Demonstrates the high-level schematic of our proposed approach - Iterative Decoding of AI Agents
+Figure 1: General pipeline for inference-alignment approaches for AI agents.
+AI agents demonstrate remarkable potential across a wide range of applications, yet they still face challenges in complex tasks requiring strategic planning and multimodal reasoning.
+For example, on challenging benchmarks such as Sketch2Code (Li et al., 2024b), Text2SQL (Li et al., 2024a), Intercode (Yang et al., 2023a), and Webshop (Yao et al., 2023), even state-of-the-art agents can only achieve 20–30% accuracy.
+While post-training via SFT and RL (Ouyang et al., 2022a, b; Bai et al., 2022) has been an effective tool to enhance the capabilities of generative models; however, it is not directly applicable to AI agents.
+This is because AI agents frequently operate as black-box systems with no access to the internals (Liu et al., 2024; Fu et al., 2024).
+Hence, we focus on inference-time approaches which are (i) API-friendly (applicable to closed-source models accessible via commercial APIs) and (ii) compatible with various underlying models.
+At inference time, alignment strategies broadly iterate over three components: sampling, evaluation, and feedback as illustrated in Figure 1.
+A typical inference-time iteration involves sampling one or more outcomes from the generative model, evaluating them using a judge/reward, and then generating feedback to improve subsequent outcomes.
+The role of sampling and evaluation is well-studied in literature with efficient and scalable algorithms based on best-of-N (BoN) approaches (Nakano et al., 2021; Beirami et al., 2024).
+BoN-based approaches operate by sampling multiple responses and selecting the best according to a verifier, effectively leveraging the sampling and evaluation steps.
+While BoN and variants (Nakano et al., 2021; Beirami et al., 2024; Verdun et al., 2025) are highly parallelizable, allowing responses to be sampled and evaluated concurrently to reduce wall-clock latency, it may not be compute-optimal.
+That is, to achieve a desirable outcome, BoN may require generating a large number of samples and evaluating each independently.
+This parallelism trades off compute efficiency for lower latency.
+Therefore, when compute is constrained, simply generating more samples is not a viable path to better performance, highlighting the need for a more adaptive approach.
+More importantly, a key limitation of BoN approaches is their inability to incorporate feedback mechanisms for iterative refinement.
+To address the above-mentioned limitations, recently, several sequential approaches have emerged (Madaan et al., 2023; Yang et al., 2023b; Chen et al., 2023; Robeyns et al., 2025; Jiang et al., 2024) that aim to explicitly integrate feedback for iterative refinement.
+This trend towards self-correction is evident in techniques such as self-refine (Madaan et al., 2023), where models iteratively improve their outputs based on self-critique, and try-again strategies (Yang et al., 2023b) that prompt the LLM to generate alternative responses when initial attempts fail.
+Other approaches include (Chen et al., 2023), where LLMs identify and correct errors in their own generated code.
+Additionally, recent advancements in agentic systems have showcased their ability to self-improve (Robeyns et al., 2025; Jiang et al., 2024).
+Figure 2: Overview of Iterative Agent Decoding (IAD), highlighting the sampling, evaluation, and feedback components of inference-time alignment.
+The figure illustrates how feedback is constructed from verification signals and used to iteratively refine model outputs.
+While existing methods creatively use feedback, they do not systematically study it.
+They often lack a focused analysis of how feedback should be designed, integrated, or optimized.
+While prior approaches (Chen et al., 2023; Yang et al., 2023b; Madaan et al., 2023) demonstrate the utility of adaptive learning with feedback, a clear understanding of the mechanisms by which LLMs learn from this feedback, especially with scalar feedback or score, is still developing.
+As a result, the role of feedback in inference-time alignment, especially in black-box agentic environments, remains under-examined.
+We aim to address this gap by focusing on feedback and systematically understanding its impact on test-time scaling of agentic workflows.
+Specifically, we focus on inference-time alignment strategies designed to maximize performance under a fixed compute budget, where compute is defined concretely as the total number of API calls or forward passes through the model.
+In such compute-constrained settings, incorporating feedback becomes critical for effective test-time scaling.
+Thus, while BoN strikes a balance between latency and compute, approaches that leverage feedback are better suited when compute is the principal focus.
+Our contributions (Understanding the role of feedback).
+To understand the role and impact of feedback on inference-time alignment, we require a unified framework that flexibly integrates diverse forms of feedback, guided by a verifier, without relying heavily on prompt engineering, which can be costly and fragile.
+To this end, we introduce Iterative Agent Decoding (IAD), a sequential framework designed to incorporate various forms of feedback to understand the role of feedback in inference-time alignment of agents.
+Through IAD we uncover several intriguing results, detailed below.
 1.
-An iterative decoding framework (IAD): We propose Iterative Agent Decoding (IAD), an efficient iterative framework where responses are refined through verifier-guided feedback, enabling improved performance in black box structured generation tasks.
-IAD is designed to extract maximal information from available feedback, including sparse or noisy scores.
+Accuracy vs compute: budget-constrained scaling with feedback: We demonstrate that feedback plays a critical role in budget-constrained settings and helps achieve up to a 10% accuracy gain over strong feedback-free baselines with the same budget.
+However, this margin narrows with increased budget, which allows for the generation of more candidates (e.g., larger $N$ in BoN) or using more capable models like Gemini-1.5 to Gemini-2.0 and Gemini-2.5 (cf.
+Figure 3).
 2.
-Empirical evaluation across agentic tasks: We evaluate IAD on three complex structured output tasks, demonstrating consistent improvements over strong baselines, while maintaining efficiency.
-IAD minimizes reliance on crafted exemplars and leverages feedback-driven reconditioning, making it task-agnostic and model-agnostic.
+Impact of adaptive feedback beyond sampling gain: We perform controlled experiments to capture the gains from feedback beyond the diversity gains from sampling, and demonstrate that the gains in IAD arise from adaptive, feedback-guided refinement, yielding performance improvements of 4–8% in both Sketch2Code and Text2SQL.
+This also highlights that feedback plays a crucial role in agentic tasks and scenarios where sampling diversity is low.
 3.
-Disentangling feedback gains in IAD from sampling: We conduct controlled experiment to isolate the effect of adaptive feedback in IAD from sampling diversity under low-temperature sampling, where IAD continues to improve, while BON performance stagnates, confirming that IAD’s gains are driven by adaptive, feedback-guided refinement, not sampling diversity alone.
+Design and role of feedback form: We investigate the impact of different forms of feedback, primarily scalar and textual on inference-time alignment.
+Textual feedback is relatively straightforward to integrate, as it can be directly incorporated into the model prompt.
+However, designing effective feedback from scalar rewards and preferences remains underexplored.
+In IAD, we focus on methods to extract useful signals from scalar rewards and demonstrate that the performance gains crucially depend on the design of proper feedback.
+Our experiments show that transforming scalar feedback into directional prompts yields gains of up to 6–7% over baselines in Sketch2Code and Text2SQL.
 4.
-Inference scaling behavior: We highlight the inference-time scaling of performance improvements where we show that both IAD and BON improve with increased test-time compute when guided by an optimal verifier.
-However, IAD benefits more significantly from high-quality feedback, leading to greater performance gains.
-Additionally, as model capability increases, BON performance improves, narrowing the gap with IAD.
-5.
-Importance of verifier quality for inference scaling: We show that high-quality verifier feedback is critical for inference scaling.
-Self-evaluation or sub-optimal verifiers often fail in complex agentic scenarios, whereas optimal verifier-guided alignment enables consistent refinement.
-6.
-Robustness under sparse and noisy feedback: To further illustrate the impact of verifier quality, we conduct controlled experiments by sparsifying and adding noise to the verifier signal.
-Results show that IAD remains robust under mild sparsity and noise, but performance for both IAD and BON degrades under extreme conditions, due to the lack of reliable feedback.
-Remarks: Several multi-turn/multi-attempt iterative approaches focus on train-time methods (Qu et al., 2024; Zhou et al., 2024; Shani et al., 2024), making them unsuitable for black box inference.
-Inference-time methods (Madaan et al., 2023; Chao et al., 2024; Mehrabi et al., 2024; Muennighoff et al., 2025) have explored iterative refinement in specific contexts like jailbreak, math, and QA, but primarily for simpler tasks rather than complex agentic settings requiring structured generations like HTML and SQL.
-Specifically, Chao et al.
-(2024); Mehrabi et al.
-(2024) iteratively refine prompts for adversarial attacks.
-Meanwhile, Madaan et al.
-(2023) and its variants emphasize on LLM self-evaluation rather than focusing on the design choices for iterative refinement based approach for structured output generation.
-2 AI Agent Alignment
-In this section, we formalize the problem of inference-time alignment for AI agents operating in black box settings, particularly in complex structured-generation tasks such as HTML and SQL synthesis in multi-modal environments like Sketch2Code, Text2SQL, and Webshop.
-Given a task distribution $\mathcal{T}\in\Delta$, we define the optimal but inaccessible agent policy $\pi^{*}(\cdot|x)$, which generates the ideal structured response $y^{*}\sim\pi^{*}(\cdot|x)$ for a given input prompt $x$.
-However, in the black box setting, we only have access to a reference policy $\pi_{0}(\cdot|x)$, which may lack some capabilities needed to complete the task.
-Figure 3: Sketch2code: This figure provides a comparison of IAD (ours) against Best-of-N sampling (SoTA) and single-turn generation with Gemini-1.5-Pro with temperature =0.6 w.r.t metrics - (a) Layout Similarity (b) TextIoU (c) ImageIoU across varying the number of generations (N).
-Figure demonstrates IAD outperforms BON consistently across N, but as N increases the gap reduces.
-Note that (b) and (c) are not directly optimized for this plot and test the generalization performance.
-This demonstrates the consistency in performance across models and metrics
-The core challenge is to optimize inference-time decisions such that the generated response from $\pi_{0}(\cdot|x)$ is better aligned with the outputs of $\pi^{*}(\cdot|x)$.
-To formally quantify this alignment gap, we define the inference-time misalignment as an $f$-divergence between the optimal and reference policies as $d_{f}(\pi^{*}(\cdot|x),\pi_{0}(\cdot|x))$, where $d_{f}(\cdot,\cdot)$ measures how far the black box agent deviates from the ideal structured output distribution.
-We explore methods to reduce this divergence using iterative inference-time strategies, which we detail in the following sections.
-We assume access to a verifier function $R(x,y)$, which evaluates the quality of a response $y$ given the prompt $x$ (further details in Appendix).
-Task specific details: In Sketch2code, the reference policy $\pi_{0}(\cdot|x)$ generates structured HTML $y$ given the prompt $x=[s,c]$ with wireframe sketch $s$, noisy text prompt $c$.
-The verifier $R(x,y)$ evaluates alignment with the target layout, textual coherence, and semantic correctness, ensuring that the generated HTML accurately reflects the sketch’s structure.
-In Text2SQL, $\pi_{0}(\cdot|x)$ maps a natural language query $c$ to an SQL query $y$ conditioned on the database schema $x$.
-The verifier $R(x,y)$ assesses syntactic validity and execution correctness, identifying errors in joins, logical consistency, and query structure.
-Finally, in WebShop, the agent’s actions are conditioned on the full interaction history up to the current time step, i.e $x_{t}=[x,y_{<t}]$, where $x$ is the initial user query and $y_{<t}$ represents all previously generated product page interactions.
-Based on this history, the reference policy $\pi_{0}(\cdot|x_{t})$ generates the next action $a_{t}$ which involves searching or clicking and the corresponding product description.
-Model Layout. Txt IoU Img IoU
-Single-Turn Approaches
-InternVL2-8b 4.01 4.89 1.41
-Llava-1.6-8b 8.01 9.26 1.95
-Claude-3-Sonnet 14.22 15.85 6.62
-GPT-4o-Mini 16.29 20.84 0.72
-Claude-3-Opus 17.11 18.09 8.32
-Claude-3-Haiku 17.52 20.60 2.72
-Gemini-1.5-Flash 17.9 17.50 10.77
-Gemini-1.5-Pro 18.25 18.20 12.69
-GPT-4o 19.20 17.12 16.19
-Gemini-1.5-Flash (CoT) 19.84 19.13 10.02
-Claude-3.5-Sonnet 22.26 25.33 9.21
-Multi-Turn Approaches (Gemini-1.5-Flash)
-Sk2code (N=2)∗∗ 19.41 20.45 11.81
-Self-Refine (N=2) 19.51 19.35 10.71
-BON (N=2) 21.45 20.1 13.5
-IAD (N=2) 24.78 23.01 15.29
-IAD-fb (N=2, K=2)∗∗ 24.66 23.4 14.69
-Self-Refine (N=4) 19.97 19.11 11.74
-Sk2code (N=4)∗∗ 20.41 21.46 12.67
-BON (N=4) 24.02 22.59 15.91
-IAD (N=4) 25.97 24.13 16.98
-IAD-fb (N=4, K=4)∗∗ 26.61 24.62 17.36
-Self-Refine (N=6) 19.89 18.91 11.61
-Sk2code (N=6)∗∗ 21.43 21.53 13.78
-BON (N=6) 25.75 22.91 17.67
-IAD (N=6) 26.75 24.91 19.12
-IAD-fb (N=6, K=6)∗∗ 27.95 24.99 19.01
-Table 1: Sketch2Code: Performance comparison between single-turn and multi-response generation approaches.
-For each of the multi-response generation method Layout score acts as the reference metric (temperature =0.6).
-Table demonstrate that IAD (Ours) consistently outperform SoTA baseline by >3-4% margin (absolute).
-∗∗ indicates those approaches leverage LLM-judge (oracle) feedback, although the LLM judges are similar but not exactly the same.
-Numbers taken from (Li et al., 2024b).
-N denotes the number of LLM calls for generating the HTML (>2000 tokens) and K represents the calls to LLM judge for getting feedback (<200 tokens).
-We have kept the judge-prompt similar to Sk2code (Li et al., 2024b)
-2.1 Limitation of Prior approaches
-In this section, we first provide a brief description of the baseline approaches and then discuss their pros and cons in this context.
-Single-turn Approaches: In single-turn approaches, the response $y\sim\pi_{0}(\cdot|x)$ is directly generated from the reference agent policy, appending some instructions to accomplish the task.
-This method is straightforward, fast, and does not rely on a verifier, making it applicable even in verifier-agnostic settings.
-However, as evident from Figure 1, direct generation — even with models like Gemini-1.5-Pro, Gemini-1.5-Flash, GPT-4, and Claude 3.X — remains highly sub-optimal for complex tasks like Sketch2Code and Text2SQL, also highlighting the difficulty of these tasks.
-Thus in single-turn generation, the performance is limited by the quality of the reference policy $\pi_{0}(\cdot|x))$ where larger f-divergence indicates greater misalignment.
-BON sampling: Best-of-N sampling improves upon single-turn generation by drawing $N$ i.i.d samples from the reference policy $\pi_{0}(\cdot|x)$ and selecting the highest-reward response based on the verifier $R(x,y)$.
-BON is simple, parallelizable, and computationally efficient and doesn’t rely on logits/model access thus applicable to black box agentic scenarios.
-It works even with scalar rewards and has been shown to achieve near-optimal trade-offs between win rate and KL divergence (Yang et al., 2024b; Beirami et al., 2024; Gui et al., 2025).
-Despite its advantages, BON remains limited by the quality of the reference policy lacks the ability to iteratively refine responses based on verifier feedback.
-For example: BON cannot incorporate targeted feedback, such as refining specific HTML structures in Sketch2Code or correcting systematic SQL errors in Text2SQL (further details in exp section)
-Figure 4: Sketch2code: This figure provides a comparison of IAD with Best-of-N sampling (BON) and single-turn generation with Gemini-1.5-Flash with temperature =0.6 for the metrics - (a, d) Layout Similarity (b, e) TextIoU (c, f) ImageIoU across varying the number of generations (N).
-Top 3 rows, the optimization is done taking Text IOU as the verifier and the bottom 3 rows with Image IOU as the verifier.
-So, this also shows both the generalisability and performance improvement of IAD over baselines.
-3 Proposed Methodology
-To address the limitations of the existing approaches, as discussed in the previous section, we propose an iterative procedure for inference time alignment of AI agents in this section.
-Our key insight: We begin by emphasizing that any iterative optimization approach generally follows three basic steps: (1) a generator to produce responses, (2) a verifier or reward function to evaluate them, and (3) a feedback mechanism to guide the next generation.
-This high-level loop is standard across any iterative frameworks, and novelty typically arises in how these components—particularly the feedback mechanism—are instantiated and integrated.
-Mathematically, this implies that if we can design a feedback mechanism over multiple iterations, denoted as $f(y_{1},y_{2},\dots,y_{T})$ , such that the such that the divergence $d_{f}(\pi^{*}(\cdot|x),\pi_{0}(\cdot|x,f(y_{1},y_{2},\dots,y_{T})))$ reduces, then we can say that feedback mechanism is effective.
-progressively decreases, then the feedback mechanism can be considered effective.
-In other words, the iterative process is successfully guiding the responses towards optimality.
-Motivated by this, we introduce our proposed approach: Iterative Agent Decoding (IAD).
-Proposed approach:
-Figure 5: Sketch2code: (a) Compares the performance of IAD at low temperature across varying number of generations to disentangle the effect of stochasticity and the improvement from iterative feedback based on IAD.(b) Comparison with BON at temperature = 0.1, which shows BON improvement ceases after 2 iterations Both (a, b) highlights the improvement due to adaptive refinement with verifier(c) Illustrates the performance of our proposed approach over 4-turns w.r.t Layout similarity score for 6 examples (temperature = 0.6).
-We propose an iterative refinement approach for black box AI agents, where response generation is guided by a verifier to progressively improve outputs.
-Step 1: At each step $t$, we sample a candidate response with the reference policy $\pi_{0}(\cdot|x)$ and iteratively refine its outputs.
-$y_{t+1}\sim\pi_{0}(\cdot|x,\hat{y}_{t},s_{t},p_{t}),$ (1)
-where $\hat{y}_{t}$ is the best response from previous iterations, and $p_{t}$ encodes prompt-based guiding instructions for ex: Surpass the best response, avoiding previous mistakes.
-This represents the more simplistic setting.
-Specifically in our algorithm, we condition on the best response $\hat{y}_{t}^{b}$ and worst response $\hat{y}_{t}^{w}$ over the iterations along with the basic instruction prompt to provide clear signals.
-Step 2: Verifier-guided selection Among the generated response and the previous best response, we select the one maximizing the reward function:
-$\hat{y}_{t+1}=\arg\max_{y\in(y_{t},\hat{y}_{t})}R(x,y)$ (2)
-Our method also has the flexibility to incorporate critique-based feedback, such as LLM-judge, which identifies specific areas needing improvement.
-More specifically we keep a track of both the best response $\hat{y}_{t+1}^{b}=\arg\max_{y\in(y_{t},\hat{y}_{t}^{b})}R(x,y)$ and similarly the worst response $\hat{y}_{t+1}^{w}=\arg\min_{y\in(y_{t},\hat{y}_{t}^{w})}R(x,y)$.
-Thus this extends to
-$y_{t+1}\sim\pi_{0}(\cdot|x,\hat{y}_{t}^{b},\hat{y}_{t}^{w},p_{t}fb_{t})$ (3)
-where $fb_{t}$ highlights specific components (e.g., incorrect tags in HTML, invalid SQL joins), guiding refinement at a more granular level in an iterative fashion.
-3. Acceptance Criterion : A new response $y_{t+1}$ is accepted if it provides an improvement over the previous best, $R(x,y_{t+1})-R(x,\hat{y}_{t})>0$. T
-4.
-Iterative Refinement: The accepted response $\hat{y}_{t}$ updates the context for subsequent generations, progressively re-weighting the proposal distribution toward higher-quality outputs.
-By conditioning on $\hat{y}_{t}$, the sampling process is guided towards responses with higher rewards, reducing the performance gap via achieving higher rewards, as shown in experimental results.
-Figure 6: Sketch2code: (a)Represents the performance of IAD with N = 2 w.r.t Layout score with varied sparsity NS (No Sparsity), LS (Low Sparsity) HS (High Sparsity) and ES (Extreme) (b) Performance comparison of IAD with BON with N = 2 at varied sparsity levels which shows the gains over BON reduces with high sparsity (c) Performance comparison of IAD with BON with N = 2 at varied noise levels of reward score which shows gains reduces over BON.
-However, IAD still improves over baselines under Sparse and Noisy feedback
-Remark (Why it works?).
-The generation of better responses occurs through stochastic sampling in each iteration from the base model, conditioned on the best response so far (and the worst candidate), followed by a pairwise comparison from the verifier.
-This feedback mechanism helps guide the generator to sample responses with higher expected rewards over the iterations.
-Practically, we achieve this by incorporating prompts like “Improve upon the best response while avoiding mistakes from the worst response.” Additionally, explicit feedback from a judge (e.g., verifier critiques or an LLM acting as a judge) accelerates the improvement process by providing targeted guidance.
-Intuitively, this approach is analogous to zeroth-order optimization (Jamieson et al., 2012; Yang et al., 2024a), where two sampled values from the objective function are sufficient to guide the optimization process toward the maximum.
-Similarly, feedback on the best and worst responses helps steer the model toward generating responses that maximize rewards, reinforced through system prompts.
-(Further details in Appendix, Section B.3)
-Figure 7: Text2SQL: (a)Accuracy comparison between the best-of-N method and our approach for Gemini-1.5-flash.
+Sensitivity to feedback quality: We study the sensitivity and robustness of inference alignment to feedback quality via controlled sparsity and noise in feedback signals.
+We see that IAD remains effective under moderate degradation, but its performance declines with increasing noise.
+Under high sparsity or heavy noise, the performance gains over sampling-based methods drop by approximately 4–5%, highlighting the importance of feedback fidelity for reliable test-time scaling.
+2 Inference-time Agent Alignment
+In the black box agentic settings, we only have access to a reference policy $\pi_{0}(\cdot|x)$, whose internals are inaccessible.
+The core challenge is to optimize inference-time decisions such that the generated response from $\pi_{0}(\cdot|x)$ is better aligned with the outputs of optimal policy $\pi^{*}(\cdot|x)$.
+We assume access to a verifier function $R(x,y)$, which evaluates the quality of a response $y$ (further details in Appendix).
+Figure 3: Test-time Scaling with Feedback in Sketch2code: This figure provides an accuracy vs compute performance comparison of IAD (ours) against Best-of-N sampling (SoTA) and single-turn generation with Gemini-1.5-Pro w.r.t metrics - (a) Layout Similarity (b) TextIoU (c) ImageIoU across varying the number of generations (N).
+The figure demonstrates IAD outperforms BoN consistently across N, but as N increases, the gap reduces.
+Fig (d, e, f) provides a comparison with improved model capabilities from Gemini-2.0-Flash, Gemini-2.5-Flash, Gemini-2.5-Pro respectively, shows that the benefit of feedback-based refinement diminishes with more capable models.
+2.1 Iterative Agent Decoding
+To study the role of feedback, we introduce Iterative Agent Decoding (IAD), a general sequential framework designed to incorporate various feedback forms for understanding the role of feedback in inference-time alignment of agents.
+Figure 4: Test-time Scaling with Feedback in Text2SQL: (a)Accuracy vs compute comparison between the best-of-N method and our approach for Gemini-1.5-flash.
 (b) Pass@K performance comparison between queries generated using our approach and Few-shot ICL with Gemini-1.5-flash.
@@
 (f) Majority@K performance comparison between queries generated using our approach and Few-shot ICL with Gemini-1.5-pro.
-Key Novel Aspect of IAD : While IAD shares a high-level structure with prior approaches including Self-Refine (Madaan et al., 2023) and its variants (Chen et al., 2023; Yang et al., 2023), it fundamentally differs in the design and integration of the feedback mechanism.
-Prior methods like Self-Debug and ReAct use binary or execution-based signals (e.g., retry if the output fails).
-This is limiting when feedback is sparse, noisy or especially scenarios with sub-optimal LLM verification.
-These methods typically use the score as a pass/fail indicator without leveraging the full informational content of the feedback which might restricting their ability to guide clear improvements in complex tasks.
-In contrast, IAD is designed to extract and utilize as much signal as possible from the available feedback.
-Rather than relying solely on absolute reward scores, IAD transforms score-based or comparative feedback into structured guidance by identifying the best and worst responses at each iteration.
-These are explicitly fed back into the model through prompt conditioning, providing a clear directional signal for improvement.
-Highlighting the Impact of Adaptive Feedback in IAD Beyond Sampling Gains Many prior works (Chen et al., 2023; Madaan et al., 2023) lack a clear experimental setup to disentangle the true source of improvement in their approaches.
-A key concern thus exists is whether the gains arise from stochastic sampling (as in Best-of-N strategies like BON), or from the actual effectiveness of self-LLM feedback.
-However, most prior methods do not include comparisons against stochastic sampling baselines like BON, leaving it unclear whether their feedback mechanisms meaningfully contribute to performance improvements.
-To address this, we conduct controlled experiment to isolate the effect of adaptive feedback in IAD from sampling diversity.
-Specifically, we reduce the generation temperature (to 0.1, 0.05, and 0.0), thereby minimizing randomness in the generation process.
-We then evaluate the performance of both IAD and BON over multiple iterations under these low-stochasticity settings as in Figures 5, 9
-Key Takeaway IAD consistently improves across iterations even with minimal sampling noise, whereas BoN performance stagnates.
-This demonstrates that the gains from IAD are not merely due to diversity in generation, but rather from the verifier-guided adaptive feedback-driven refinement (further details in Section 4.1)
-Figure 8: Performance comparison of IAD, BON with improved Model- Gemini-2.0-Flash on (a) Text2SQL and (b) Sketch2Code across varying iterations.
-Improvement in Multi-modal understanding of Gemini-2.0 over Gemini-1.5 as evident in performance (Baseline improves from 18 to 23).
-In Text2SQL, performance remains comparable to Gemini-1.5-Pro.
-In Text2SQL, performance of BON improves and gap reduces although still IAD outperforms.
-4 Experimental Analysis
-In this section, we provide detailed discussion on the experimental analysis with respect to the agentic tasks below.
-For all our experiments, we have utilized NVIDIA A100-SXM4-40GB GPU with 40GB of VRAM, running on CUDA 12.4 and driver version 550.90.07.
-4.1 Sketch2code Results and Analysis
-Evaluation metrics and Baselines.
-We evaluate our approach against SoTA baselines using (i) Layout Similarity, (ii) Visual IoU, and (iii) Text IoU with reference HTML (Li et al., 2024b).
-These metrics strongly correlate ($\approx 90\%$) with human satisfaction (Li et al., 2024b) (details in Appendix).
-We use layout similarity as a verifier alongside LLM-as-judge (Li et al., 2024b) to refine the generations for both BON (Beirami et al., 2024) and IAD.
-We compare our proposed approach IAD against single-turn SoTA models (GPT-4o, Claude-3, InternVL2, Gemini-1.5-Flash, CoT) and multi-response approaches (BON, Sk2code).
-We compare the performance on the same evaluation setup and dataset as used in (Li et al., 2024b).
-Initial results.
-Figure 1 and Table 1 qualitatively and quantitatively show that single-turn approaches, including SoTA models, fail to maintain the layout structure, block positions, text placement, and sizing, leading to poor alignment with the reference layout.
-These models score low across all three metrics.
-Multi-response methods significantly improve performance, even with N=2 samples.
-BON, using the weaker Gemini-1.5-Flash, outperforms single-turn models.
-With N=4, BON surpasses SoTA single-turn models by $2-4\%$ (absolute improvement), accurately capturing block positioning, title blocks, and overall layout.
-Performance monotonically improves with more responses, with layout scores rising from $20.41$ to $25.7$ as N increases to six.
-Benefits of IAD (ours).
-BON struggles with fine-grained layout details and often repeats positional errors across all N generations (Figure 1).
-In contrast, IAD iteratively refines responses, leading to a $\approx 3-4\%$ improvement (absolute) over BON and single-turn SoTA (Claude) in just two iterations (N=2), even with the simpler Gemini-1.5-Flash model (Table 1).
-At each iteration, IAD incorporates the best and worst HTML outputs as context, along with refinement instructions (with and without LLM judge).
-This enables progressive learning of layout components and image semantics, significantly enhancing output quality.
-As iterations increase, IAD achieves a high layout score of 26.75, outperforming all baselines with the same number of generations.
+The IAD process begins with an initial input $x$ and iteratively improves a candidate solution $\hat{y}$ over a series of iterations.
+Each iteration $t$ is composed of the following interconnected stages:
+Step 1: Sampling and Generation.
+At each step $t$, a new candidate solution $y_{t+1}$ is sampled from the reference policy $\pi_{0}$.
+This policy is conditioned on the input $x$ along with the best solution found in previous iterations $\hat{y}_{t}$ and the feedback $fb_{t}$
+$y_{t+1}\sim\pi_{0}(\cdot|x,\hat{y}_{t},s_{t},p_{t}),$ (1)
+Feedback $fb_{t}$ encodes guiding instructions incorporated into the model’s input prompt for ex: Surpass the best response, avoiding previous mistakes or Improve upon the previous best solution by addressing the following errors…
+Step 2: Evaluation and Selection : The newly generated candidate $y_{t+1}$ is then evaluated against the current best $\hat{y}_{t}$ using the reward function or verifier $R(x,y)$.
+The better solution is selected and carried forward to the next iteration as the new best-so-far
+$\hat{y}_{t+1}=\arg\max_{y\in(y_{t},\hat{y}_{t})}R(x,y).$ (2)
+IAD has the flexibility to incorporate critique-based feedback, such as LLM-judge, which identifies specific areas needing improvement.
+Step 3: Feedback Integration.
+One of the most crucial components in iterative decoding approaches is its ability to incorporate rich, structured feedback to guide the next generation step.
+The evaluation from Step 2 can be either:
+(i) textual feedback, which provides natural language instructions or critiques;
+(ii) scalar feedback, such as numerical scores or preference comparisons.
+Textual feedback is relatively easy to integrate, as it can be directly injected into the model’s context or prompt.
+However, generating high-quality textual feedback is often expensive, requiring either powerful LLM judges or human annotation, and may vary in precision and consistency.
+Challenge in Feedback (Score to Text) Extraction and Our Key Approach: On the other hand, scalar feedback is significantly easier to obtain; however, translating these signals into actionable guidance remains a key challenge.
+Directly appending the scalar value to the prompt, as done in several prior works (Yang et al., 2024), does not leverage the full informational content of the signal and often fails to yield meaningful improvements.
+Hence, we design a feedback integration method in IAD that translates scalar scores into structured directional feedback.
+At each iteration, IAD transforms score-based or comparative feedback into structured guidance by identifying the best and worst responses at each iteration:
+$y^{t}_{\text{best}}$ $=\arg\max_{y\in{y_{t+1},\hat{y}_{t}}}R(x,y),$ (3)
+$y^{t}_{\text{worst}}$ $=\arg\min_{y\in{y_{t+1},\hat{y}_{t}}}R(x,y)$ (4)
+These are explicitly fed back into the model through prompt conditioning. The feedback is constructed by concatenating the instruction with the best and worst responses:
+$fb_{t}=\texttt{[Instruction]}\oplus\texttt{Best:}y^{t}_{\text{best}}\oplus\texttt{Worst:}y^{t}_{\text{worst}}$ (5)
+where the Instruction states generic guidance like Improve upon the best response while avoiding mistakes from the worst. This feedback thus provides a clear directional signal for improvement.
+The accepted response $\hat{y}_{t}$ updates the context for subsequent generations, progressively re-weighting the proposal distribution toward higher-quality outputs.
+By conditioning on $\hat{y}_{t}$, the sampling process is guided towards responses with higher rewards, reducing the gap between the reference distribution $\pi_{0}(\cdot|x,\hat{y}_{t})$ and the optimal distribution $\pi^{*}(\cdot|x)$, as observed in experimental results.
+Example: For instance, in an SQL code generation task, the model might initially produce a non-functional or erroneous SQL statement.
+However, in the next iteration, it generates a different SQL code, and through verifier comparisons, we determine which version is better—e.g., if the later version passes more test cases, we infer that the model should move in that direction.
+Using an LLM as a judge makes this refinement process more targeted, as it can provide explicit feedback on errors like "This condition is incorrect", "Fix the syntax here", or "This table join is unnecessary" and suggest improvements at each iteration.
+Figure 5: Sketch2code: This figure provides a comparison of IAD (ours) with Best-of-N sampling (SoTA) and single-turn generation with Gemini-1.5-Flash for the metrics - (a, d) Layout Similarity (b, e) TextIoU (c, f) ImageIoU across varying the number of generations (N).
+Top 3 rows, the optimization is done taking Text IOU as the verifier and the bottom 3 rows with Image IOU as the verifier.
+So, this also shows both the generalisability and performance improvement of IAD over baselines.
+Figure 6: Beyond Sampling in Sketch2code: (a) Compares the performance of IAD at low temperature across varying number of generations to disentangle the effect of stochasticity and the improvement from iterative feedback based on IAD.
+(b) Text2SQL :Compares the accuracy of IAD on BIRD development set at low temperature across varying number of generations to highlight the improvement from iterative feedback based on IAD.
+Baseline is Gemini-1.5-pro.
+BON remains similar as stochasticity improves across the number of responses.
+Figure 7: Sketch2code: (a) Represents the performance of IAD with N = 2 w.r.t Layout score with varied sparsity of feedback which shows gains reduces over BON.
+(b) Performance comparison of IAD with BON with N = 2 at varied noise levels of feedback, which shows gains reduced over BON.
+However, IAD still improves over baselines under Sparse and Noisy feedback.
+Figure (c) Represents the performance of Self-Refine (Madaan et al., 2023) based approaches across iterations using LLM as a judge.
+(d) Represents the performance of IAD with a scalar score, which shows monotonic improvements, highlighting the importance of the optimality of feedback and verification.
+3 Experiment Setup
+In this section, we provide a detailed discussion on the experimental analysis with respect to the agentic tasks below.
+For all our experiments, we have utilized an NVIDIA A100-SXM4-40GB GPU with 40GB of VRAM, running on CUDA 12.4 and driver version 550.90.07.
+To enable a comprehensive evaluation, we organize the experimental setup into three components: (1) the benchmarks used to assess agentic capabilities across diverse modalities, (2) the base models serving as generative backbones, and (3) the set of baselines for comparative analysis.
+3.1 Benchmarks
+We perform empirical analysis on 4 challenging agentic environment benchmarks detailed as follows.
+Text2SQL tasks map natural language questions to executable SQL queries over structured databases.
+These tasks require deep reasoning to interpret user intent and generate syntactically and semantically correct queries.
+We use the BIRD benchmark Li et al.
+(2024a), which includes 12,751 question-SQL pairs spanning 95 databases and 37 professional domains.
+Performance is evaluated using execution accuracy (EX), where a prediction is correct if it yields the same results as the ground truth query when executed.
+Sketch2code (Li et al., 2024b) tests the multi-modal abilities of agents by converting wireframe sketches into functional HTML prototypes.
+The task requires aligning visual cues with structured code, handling layout ambiguities, and generating precise UI structures.
+Evaluation is based on three metrics: Layout Similarity (IoU over UI components), Text IoU (text alignment accuracy) and Image IoU (CLIP-based visual similarity).
+These metrics strongly correlate with human judgment.
+Intercode (Yang et al., 2023a) evaluates agents in structured programming tasks across four domains: Bash, Python, SQL, and CTF.
+Each task involves a series of agent decisions, including command execution and error handling, requiring both syntactic precision and logical planning.
+Performance is measured using reward signals derived from task execution correctness.
+We focus on the Bash domain for our experiments, where agents must generate correct shell commands based on natural language instructions.
+Webshop (Yao et al., 2023) simulates a real-world online shopping environment, requiring agents to perform long-horizon, language-guided decision-making.
+Given product queries, agents interact with webpages (search, click, select) to find matching items.
+Key evaluation metrics include: Success Rate (SR): whether the selected item satisfies all constraints.
+Progress Rate (PR): how closely the agent’s actions align with task goals.
+3.2 Base Models
+We experiment with a wide range of large language and vision-language foundation models across all four benchmark tasks to study the generality of IAD across different agentic settings.
+Here we outline the base models used for each benchmark.
+For Sketch2Code, we use Gemini-2.5-Pro (Google DeepMind, 2024), Gemini-2.0-Flash, Gemini-1.5-Pro, and Gemini-1.5-Flash (Team et al., 2024) as base models to compare test-time scaling approaches with and without feedback.
+Furthermore, we include InternVL2-8b (Chen et al., 2025), Llava-1.6-8b (Liu et al., 2023), Claude-3-Sonnet , GPT-4o-Mini, Claude-3-Opus, Claude-3-Haiku, Claude-3.5-Sonnet and GPT-4o (OpenAI et al., 2024) as baseline models for broader comparison.
+Vision-language models are used to handle the multi-modal input (user sketches) and generate HTML output.
+For Text2SQL, we use Gemini-1.5-Pro, Gemini-1.5-Flash, GPT-4o.
+These are used both as generative models and occasionally as judges to provide feedback.
+For Webshop, we use Gemini-1.5-Pro, Gemini-1.5-Flash, and GPT-4o to evaluate the effectiveness of IAD in long-horizon decision-making within a simulated online shopping environment.
+Additionally, we also use Lemur-70b (Xu et al., 2024), Mistral-7b (Jiang et al., 2023), Vicuna-13b-16k (Zheng et al., 2023), GPT-3.5-Turbo-16k, Text-Davinci-003 (Ye et al., 2023), DeepSeek-67b (DeepSeek-AI et al., 2024) for additional baseline comparisons.
+Finally, for Intercode (Yang et al., 2023a), GPT-4 is used as the base model for generating Bash commands and interacting in structured programming tasks.
+3.3 Baselines
+We compare and contrast IAD against a range of black-box inference-time approaches, which include: (1) Single-Turn: Zero-shot generation from the base model using only the input prompt.
+(2) Best-of-N (BoN) and its variants (Jinnai et al., 2024; Amini et al., 2024; Beirami et al., 2024; Gui et al., 2025): Sampling-based methods where multiple outputs are generated from the base model and the best is selected via a verifier.
+(3) Self-Refine (Madaan et al., 2023) and related methods such as Self-Debug (Chen et al., 2023) and Try-Again (Yang et al., 2023b): Iterative refinement approaches that use self-critique or LLM-based retry mechanisms to improve outputs.
+(4) We also compare IAD with few-shot prompting-based methods such as DIN-SQL (Pourreza and Rafiei, 2024) and DAIL-SQL (Gao et al., 2023), as well as multi-agent test-time approaches like MAC-SQL (Wang et al., 2023), to compare the performance of IAD over test-time baselines.
+We study the effect of feedback in IAD along four key dimensions: (1) Accuracy vs compute - Budget controlled scaling (2) Impact of adaptive feedback beyond sampling diversity (3) Impact of feedback modalities, (4) sensitivity to feedback quality under varying levels of noise and sparsity
+Figure 8: (a) Test-time scaling with feedback in Intercode : Comparison of IAD with baselines (GPT-4) on the accuracy–compute trade-off shows that IAD, with appropriate feedback, enables more effective test-time scaling.
+Intercode : Comparison of IAD with baselines (GPT-4) on the accuracy–compute trade-off shows that IAD, with appropriate feedback, enables more effective test-time scaling.
+(b) Test-time scaling with feedback in Webshop : Progress-rate comparison of IAD over Best of N with Self-consistency (BON-SC) and baseline on a random selection of 20% of the total evaluation set using weaker Gemini-1.5-Flash-01 (weaker model).
+Figure highlights the role of feedback in IAD with a weaker model.
+4 Experiment Results and Analysis
+4.1 Accuracy vs compute tradeoff
+We investigate how IAD scales with compute and model capacity across Sketch2Code, Text2SQL, Intercode, and Webshop.
+Our results show that with accurate feedback, IAD achieves up to 10% accuracy gains over feedback-free baselines like BoN under constrained budgets as shown in Table 3, Table 1.
+However, as compute increases, via more samples or stronger models, the performance gap narrows, indicating diminishing returns from feedback alone.
 Method Exe Acc
@@
 Table 1: Text2SQL - Execution accuracy comparison of previous works with our proposed approach
-We also evaluate Image and Text IoU scores to check for reward over-optimization.
-However, as shown in Table 1 and Figure 3, text and image similarities also improve, confirming consistency in IAD’s performance gains over BON.
-Notably, with more generations, the performance gap between IAD and BON narrows.
-We also qualitatively (appendix) and quantitatively evaluate the performance across various samples with IAD which shows improvement in scores across the iterations for examples Figure 5(c).
-The increasing trend for samples show that our approach is able to improve the performance with increased iterations/inference compute.
-To further investigate and disentangle the improvement of IAD over BON or stochastic parallel sampling, we reduce the temperature (0.0, 0.05, 0.1) and perform IAD and still observe the trend in performance as shown in 5(a) 5(b), whereas the performance of BON ceases at layout score of ~21.9 even if we increase the N.
-This clearly shows the importance of adaptive feedback to gain a higher performance in-terms of layout score of ~26.
-Figure 9: Text2SQL :Compares the accuracy of IAD on BIRD development set at low temperature across varying number of generations to highlight the improvement from iterative feedback based on IAD.
-Ablation.
-We also consider the sensitivity of the token length of the context plays a critical role in this case, where providing the entire HTML can affect the entropy of the distribution, and thus, over-conditioning can hinder structured generation by reducing diversity and exploration.
-Thus, we provide only the top 200-300 tokens of the best (and worst) HTMLs.
-However, we remark that if there were a judge to highlight which portion of the code needs to be updated, that would be more targeted.
-Hence, we incorporate LLM-judge (Gemini-1.5-Pro), which has the reference policy, and it checks with the current response and provides feedback on improvement and sometimes snippets of HTML as well (however, we restrict that to 200 tokens -5-6% of the original HTML).
-This leads to an additional improvement of  2-4% (absolute improvement) for the layout score with just two iterations and a final score of ~28 with 6 iterations, demonstrating the importance of iterative approaches with judge feedback for agent performance.
+Sketch2Code: Layout Similarity, Text IoU, and Image IoU steadily improve with more IAD iterations.
+Even with just N=2 and a weaker model (Gemini-1.5-Pro), IAD surpasses BoN and single-turn baselines by 3–4% (Figure 3, Table 3).
+But with more capable models (e.g., Gemini-2.5-Pro), BoN narrows the gap, suggesting parallel sampling benefits from model strength.
+Text2SQL: IAD outperforms few-shot CoT and BoN using the same models (Gemini-1.5-Pro/Flash), achieving higher execution accuracy with fewer LLM calls (Figure 4).
+Through just three rounds of feedback, IAD effectively corrects syntactic and semantic SQL errors, demonstrating feedback-driven compute efficiency.
+Compared against diverse baselines (DIN/DAIL/MCS/MAC-SQL) that avoid fine-tuning, IAD still leads (Table 1), reinforcing feedback’s central role.
+A similar trend is observed in Intercode Figure 8(a), which shows sequential refinement with appropriate feedback outperforms baselines with low budget.
+Webshop: IAD outperforms BoN-SC and strong baselines like GPT-4o and Gemini.
+For example, SR improves from 29.3% (Gemini-1.5-Pro) and 41.09% (BoN-SC + GPT-4o) to 44.68% (IAD + GPT-4o), a 3–4% absolute gain.
+In weaker models like Gemini-1.5-Flash, iterative refinement significantly boosts performance where BoN plateaus.
+Key Insights : We observe that feedback plays a crucial role in budget-controlled regimes—IAD consistently outperforms non feedback based baselines like BoN when compute (e.g., number of samples or model strength) is limited.
+However, as computational budget or model capability increases, the advantage of feedback-based iterative refinement diminishes.
 Figure 9: Qualitative illustration of IAD in improving the action taking capabilities of Webagent for Webshop environment (Yao et al., 2023) over direct generation, where IAD improves the search action by iteratively reviewing and refining the actions.
-We have kept the judge-prompt similar as Sk2code (Li et al., 2024b) for fairness in evaluation.
-It is important to note that IAD-fb represents that we augment the judge feedback along with our best and worst response and generic instruction prompt as described in Algorithm.
-We admit that its indeed possible to learn more specific prompts to construct a better judge via training/prompt-tuning, however that’s outside the main scope of this work.
-Overall, in all our analyses, our findings remain consistent, where IAD outperforms baselines.
-4.1.1 Text-to-SQL Results and Analysis
-In this section, we detail the experiments conducted on the BIRD text-to-SQL benchmark (Li et al., 2024a).
-Our evaluation involved two models, Gemini-1.5-pro and Gemini-1.5-flash, to test our proposed approach and compare it against baseline methods.
-IAD vs test-time-compute methods:
-We compared our method to two baselines: few-shot chain-of-thought prompting and the best-of-N approach, both implemented using the same models.
-These comparisons highlight the effectiveness of our multi-turn strategy in narrowing the gap between the reference distribution and the optimal distribution.
-Our multi-turn approach addresses common issues in single-turn generation, which can produce SQL queries with syntax and semantic errors.
-By incorporating multiple rounds of response selection and feedback, our method effectively corrects these errors.
-Moreover in the multi-turn setting, the reward model benefits from observing query execution results, allowing it to more readily identify syntax and semantic issues.
-For the few-shot prompting baseline, we employed two metrics—Pass@K and Majority@K—to evaluate the quality of the generated candidate queries.
-In order to obtain candidate queries we execute our approach multiple times with three rounds of refinement.
-As shown in Figure 7, our approach consistently produces candidate sets with higher correctness.
-In contrast, the best-of-N method aggregates results, using self-consistency; therefore, we used the original execution accuracy for its evaluation.
-To ensure a fair comparison, we computed accuracy across different numbers of LLM calls.
-In our method, we first generate a draft SQL query, followed by multiple rounds of evaluation, feedback generation, and query refinement, each contributing to the LLM call count.
-As illustrated in Figure 7, our approach consistently produce higher accuracy across different number of LLM calls compared to the best of N approach.
-Similar to the previous section, Figure 9 demonstrates the performance improvement when increasing the number of LLM calls across different temperatures.
-IAD vs Text-to-SQL baselines
-We also compare our approach with several baseline methods in the text-to-SQL domain that do not rely on fine-tuning on the BIRD train set, ensuring a fair comparison.
-These baselines originate from different families of methods, for example, MCS-SQL (Lee et al., 2024) is a variant of the best-of-N method, DIN-SQL (Pourreza and Rafiei, 2024) and DAIL-SQL (Gao et al., 2023) employ human-engineered few-shot prompts to achieve higher performance, and MAC-SQL (Wang et al., 2023) is a multi-agent approach designed specifically for text-to-SQL tasks.
-The results of these comparisons are provided in Table 2, where we achieved the highest execution accuracy by 3% margin compared to the second best method.
-4.2 Webshop Results and Analysis
+4.2 Impact of adaptive feedback
+Prior works lack a clear experimental setup to disentangle the true source of improvement in their approaches (Chao et al., 2024; Mehrabi et al., 2024; Wang et al., 2022).
+A key concern thus exists is whether the gains arise from sampling diversity or from the actual effectiveness of the feedback.
+However, most prior methods do not include comparisons against stochastic sampling baselines like BON, leaving it unclear whether their feedback mechanisms meaningfully contribute to performance improvements.
+To answer this, we conduct a controlled experiment to isolate the effect of adaptive feedback in IAD from sampling diversity.
+Specifically, we reduce the generation temperature (to 0.1, 0.05, and 0.0), thereby minimizing randomness in the generation process.
+We then evaluate the performance of both IAD and BON over multiple iterations under these low-stochasticity settings.
+Sketch2Code: BoN saturates at layout score ~21.9 even with N=6, while IAD surpasses 26 with 6 iterations.
+Figure 6 confirms that adaptive refinement—not randomness—drives gains.
+Text2SQL: Accuracy increases steadily with IAD even at low temperature settings (Figure 6).
+This validates the role of feedback in semantic correction.
+Key Insight: The results demonstrate that the gains from IAD are not merely due to diversity in generation/sampling, but rather from the verifier-guided adaptive feedback-driven refinement.
+This also highlights a critical insight that when the diversity in the policy is low, feedback based sequential approaches can perform significantly better than sampling based BON approaches.
+4.3 Design and role of feedback form
+We investigate the impact of different forms of feedback on inference alignment.
+When feedback is textual, it is relatively easy to integrate as is.
+However, designing effective feedback from scalar rewards and preferences remains underexplored.
+In IAD, we focus on various methods to extract useful signals from scalar rewards and demonstrate that performance gains crucially depend on the design of proper feedback.
+We analyze both the scenarios: 1.
+Textual feedback from LLM as a judge 2.
+Scalar rewards.
+Majority of the prior sequential approaches including (Madaan et al., 2023) have been designed with a focus on approach 1 i.e LLM as a judge and taking feedback from the same.
+However, we demonstrate that directly using off-the-shelf LLMs as judges fails to deliver consistent or monotonic improvements in complex scenarios (Figure 7).
+Our experiments reveal that performance often plateaus, fluctuates, or even degrades across iterations.
+To highlight this, we include both quantitative and qualitative examples (Appendix) where self-LLM feedback is repetitive and uninformative, failing to identify or correct meaningful issues.
+This supports our central claim: for effective inference-time optimization, near-optimal verifiers are essential, especially in complex agentic scenarios like Sketch2code, text2sql.
+How to design feedback from a scalar reward?
+A crucial challenge lies in designing meaningful and informative feedback from scalar reward or preference, since the eventual feedback needs to be in textual form.
+Thus IAD is designed to extract as much signal as possible from the available feedback.
+Rather than relying solely on absolute reward scores, IAD transforms score-based or comparative feedback into structured guidance by identifying the best and worst responses at each iteration.
+These are explicitly fed back into the model through prompt conditioning, providing a clear directional signal for improvement.
+This process not only reinforces the distinction between good and bad outputs but also enables the use of dense feedback—even from weak or indirect supervision sources—turning minimal signals into effective updates.
+Intuitively, this approach is analogous to zeroth-order optimization, where two sampled values from the objective function are sufficient to guide the optimization process toward the maximum.
+Similarly, feedback on the best and worst responses helps steer the model.
+We compare IAD with the above feedback for Sketch2code and show that it performs almost comparable with LLM judge with the reference.
 Models (PR) (SR)
@@
 For BON-SC, we generated N=3 responses, similarly we used three feedback steps for IAD.
-For all the experiments, we use greedy decoding with temperature=0.1, top p = 1.0 for our results.
-For BON-SC, we try to vary the temperature to generate more diverse responses.
-We compare IAD* against SoTA baselines, including Lemur, Mistral, Vicuna, Gemini, and GPT models, in the Webshop environment, where agents interact with webpages through search and click actions.
-Performance is measured via progress rate (PR) and success rate (SR), with SR being the key metric for exact attribute satisfaction.
-We follow the same evaluation methodology as AgentBoard (Ma et al., 2024).
-IAD* outperforms baselines, achieving  8-10% improvement (absolute) in SR across most models with Gemini-1.5 inclduing BON and achieves an improvement of over 3-4% (absolute) with very strong GPT-4o as shown in Table 3.
-To further enhance performance, we introduce feedback-based iterative refinement, where agents receive prompts like "Reflect on your previous action and improve upon it", "Improve the previous actions focusing on key-attributes", "Follow the format to be action.name-’search’ and action.params -’red tshirt….’" for allowing LLMs to iteratively refine their decisions.
-A verifier-based LLM judge determines when to iterate or finalize actions.
-We also compare IAD* with BON-SC (Best-of-N with Self-Consistency, N=3) but find that BON provides limited improvement due to low diversity in generations, even with temperature = 0.6.
-Beyond this, higher temperatures introduce irrelevant attributes, reducing effectiveness.
-Figure 11: Webshop : Progress-rate comparison of IAD over Best of N with Self-consistency (BON-SC) and baseline on random selected 20% of the total evaluation set using weaker Gemini-1.5-Flash-01 (weaker model).
-Figure highlights the importance of iterative feedback while generation.
-Evaluation was done with a weaker model and on a subset of eval examples.
-In contrast, IAD* improves search queries by refining unwanted attributes, enhancing focus on key parameters.
-Additionally, to show the effectiveness, we perform with a weaker version of Gemini-1.5-Flash-001 and perform the experiment, where we observe that it needs more iterations to actually follow the action format, without this iteration it fails to provide response in a meaningful format.
-On the other-hand, we approach the same with BON-SC which couldn’t improve much due to the issue in diverse response generation as highlighted before.
+4.4 Sensitivity to feedback quality
+We study the sensitivity and robustness of inference alignment to feedback quality via controlled sparsity and noise in feedback signals.
+Sparse Rewards.
+In many agentic tasks, dense reward signals may not be available.
+We study this setting in the context of Sketch2code (Li et al., 2024b) where we systematically sparsify our verification signal by varying the level of feedback sparsity and comparing the performance of IAD against the BON baseline with and without sparsity.
+Sparsification was achieved by providing feedback only when the verifier score exceeded a threshold; the higher the threshold, the sparser the reward signal.
+We define the sparsity levels and evaluate the performance of IAD versus BON with N=2 responses per prompt.
+We observe that as the level of sparsity increases, the performance gap between IAD and BON narrows.
+Both inference-time approaches—BON and IAD—show a decline in performance, approaching the baseline under extreme sparsity conditions.
+The primary hypothesis behind the drop in IAD performance is that IAD relies on adaptive feedback: the LLM is conditioned on the best and worst responses from previous iterations, akin to a zeroth-order optimization method.
+Feedback on both ends (positive and negative) helps steer the model toward reward-maximizing generations through updates to the system prompt.
+However, under extreme sparsity, most responses—regardless of quality—receive zero reward.
+This results in random selection and noisy update directions, limiting IAD’s ability to effectively adapt and improve.
+Noisy Verification.
+To further investigate the effect of noisy verification and how it impacts the performance of IAD compared to BON, we introduce varying levels of noise into the reward signal.
+Specifically, we add Gaussian noise with different variances to the reward scores, simulating imperfect or noisy verifier conditions.
+We then run both IAD and BON with N=2 responses per prompt and evaluate their performance under these noisy settings.
+This setup allows us to assess the robustness of IAD to reward noise and compare its stability and effectiveness relative to BON when the verification signal is noisy.
+Key Insights : We observe trends similar to those in the sparse reward setting.
+Notably, IAD remains reasonably robust to mild noise in the reward signal.
+This is because IAD relies on adaptive feedback, where the LLM is conditioned on the best and worst responses from previous iterations, effectively leveraging pairwise comparisons rather than absolute scores.
+As long as the noise is limited, the relative preference between responses is preserved.
+For instance, if one layout score is 0.5 and another is 0.3, mild noise might shift them to 0.55 and 0.24, respectively, maintaining the same ordering.
+Therefore, IAD continues to improve under reasonable noise levels.
+However, as noise increases significantly, it can flip these preferences, leading to unstable updates and a decline in performance.
 5 Conclusion
-In this work, we proposed IAD : an iterative decoding approach for AI agent alignment with black box access which highlights the effectiveness of iterative decoding (guided by a verifier) for these complex agentic tasks.
-IAD enhances prior baselines by dynamically integrating verification and critique mechanisms to iteratively refine and improve response quality.
-Experimental results across diverse domains show that our approach consistently outperforms several existing baselines including highly competitive BON highlighting its ability to bridge the gap between reference and optimal policies at inference time.
-6 Limitations
-While IAD — our proposed iterative decoding approach — improves upon prior baselines by leveraging verifier feedback, it is inherently sequential, leading to increased computational overhead compared to easily parallizable BON approaches.
-Addressing this trade-off between quality improvement and computational efficiency remains an important area for future research.
-Efficient optimization approaches, such as adaptive stopping, speculative decoding, and more efficient verifier-guided selection, could improve the efficiency in iterative decoding for agentic tasks.
-Additionally, verifier and judge plays a crucial role in our approach.
+In this work, we explore the underexplored role of feedback in inference-time alignment for black box AI agents.
+To understand the effect of feedback in inference alignment, we introduce Iterative Agent Decoding (IAD), a general sequential framework.
+Our study analyzes feedback through four lenses: (1) accuracy vs.
+compute trade-offs, (2) gains beyond sampling diversity, (3) feedback modality integration, and (4) sensitivity to feedback quality.
+Empirically, we find that feedback is especially valuable under constrained budgets, achieving up to 10% gains over feedback-free baselines.
+We also highlight the challenge of integrating diverse feedback modalities into sequential designs, which has not been critically explored in the literature.
+While textual feedback integrates naturally, representing scalar or preference signals remains an open challenge.
+We also observe that IAD’s benefits diminish under highly sparse or noisy feedback, underscoring the importance of feedback fidelity for effective alignment.
+Limitations
+While IAD, our iterative decoding approach, improves upon prior baselines by better leveraging verifier feedback, it is inherently sequential, leading to increased user-facing latency compared to easily parallelizable BoN approaches.
+Addressing this tradeoff between quality improvement, computational cost, and user facing latency remains an important area for future research, which may require properly combining these techniques with adaptive stopping, controlled decoding (Mudgal et al., 2024), speculative decoding (Leviathan et al., 2023).
+Additionally, more efficient verifier-guided selection could improve the efficiency in iterative decoding for agentic tasks.
+As we learnt, the verifier (or judge) plays a crucial role in our approach.
 Thus a more concrete investigation and selection of a judge for these challenging tasks is a valid and crucial next step of our work.
-We highlight that this work is of academic nature and has no direct or immediate harmful impacts to society.
+We highlight that this work is of an academic nature and has no direct or immediate harmful impacts on society.
 However, since this work deals with improving AI agents, it should be done under safety protocols and guidelines.
-We want to highlight that this study is limited to English language text primarily due to the nature of open-source datasets used.
+We want to highlight that this study is limited to English language text primarily due to the nature of the open-source datasets used.
 […]
@@
 Generally, a Text-to-SQL system receives a natural language question and any pertinent metadata about the tables and columns, which serves as external knowledge to aid in database comprehension.
-Consequently, such systems are responsible not only for interpreting user intent and identifying relevant information from a potentially vast set of tables and columns but also for generating SQL queries that may include multiple conditions—a process that is inherently reasoning intensive.
+Consequently, such systems are responsible not only for interpreting user intent and identifying relevant information from a potentially vast set of tables and columns but also for generating SQL queries that may include multiple conditions—a process that is inherently reasoning-intensive.
 To evaluate our proposed framework, we employ the BIRD benchmark (Li et al., 2024a), a challenging and widely used dataset in the Text-to-SQL domain.
@@
 These metrics provide a reliable way to measure the quality of the generated response and strongly correlates with human judgement.
-Evaluations are done also with LLM as judge to compare the performance.
+Evaluations are also done with LLM as a judge to compare the performance.
 3.
@@
 Figure 11: Sketch2code: Qualitative evaluation of the generated HTMLs with BON sampling (N=4) corresponding to the user-sketch (left-bottom) and reference html (left-top).
-The figures show that BON performs much better in matching the reference HTML but still misses specific properties like rectangular structue, position of text, relative positioning of smaller blocks etc.
+The figures show that BON performs much better in matching the reference HTML but still misses specific properties like rectangular structure, position of text, relative positioning of smaller blocks etc.
 Appendix B Detailed Experimental Analysis
@@
 We first compare our proposed method with the widely used few-shot prompting approach in terms of Pass@k performance and final accuracy after self-consistency (Majority@K) using execution accuracy as the metric in order to demonstrate that using our method we can generate a pool of candidates with a higher quality.
-Subsequently, we compare our approach with the best-of-N approach which is one of the strong baselines as a test-time compute approaches to demonstrate the effectiveness of the proposed framework.
+Subsequently, we compare our approach with the best-of-N approach, which is one of the strong baselines as a test-time compute approach to demonstrate the effectiveness of the proposed framework.
 Finally, we compare our method with all previously proposed test-time methods on the BIRD development set benchmark, excluding works that rely heavily on fine-tuning LLMs (Pourreza et al., 2024; Talaei et al., 2024; Maamari et al., 2024; Gao et al., 2024) for a fair comparison.
@@
 It is evident that HTMLs with higher match with the reference layout (right-top) and user sketch(right-bottom) has higher layout score and vice-versa showing that its a valid metric.
-[t] Sketch2code : Oracle Judge Prompt for providing Feedback Judge Prompt for Sketch2code (similar to (Li et al., 2024b)): Act as you are a front-end designer working with a code agent to implement an HTML webpage .
+[t] Sketch2code : Oracle Judge Prompt for providing Feedback Judge Prompt for Sketch2code (similar to (Li et al., 2024b)): Act as if you are a front-end designer working with a code agent to implement an HTML webpage.
 You are provided with two HTMLs: the first is the reference webpage, and the second one is the current implementation from the code agent.
-The task is to carefully compare the agent ’s implementation against the reference webpage , and provide feedback to help the agent make its implementation closer to the reference webpage .
+The task is to carefully compare the agent’s implementation against the reference webpage and provide feedback to help the agent make its implementation closer to the reference webpage.
 Your feedback should be specific to the differences in layouts and visual components on the two webpages.
-Don’t focus on the style components too much and focus on layout similarity and visual match with the reference webpage.
+Don’t focus on the style components too much, and focus on layout similarity and visual match with the reference webpage.
 [t] Sketch2code : Feedback from LLM as a Judge (Self Verification) Feedback provided from Self-LLM: Iter1: The HTML structure and CSS styling do not reflect the provided wireframe.
@@
 Use CSS Grid to precisely position and size all elements according to the wireframe’s proportions.
+Model Layout. Txt IoU Img IoU
+Single-Turn Approaches
+InternVL2-8b 4.01 4.89 1.41
+Llava-1.6-8b 8.01 9.26 1.95
+Claude-3-Sonnet 14.22 15.85 6.62
+GPT-4o-Mini 16.29 20.84 0.72
+Claude-3-Opus 17.11 18.09 8.32
+Claude-3-Haiku 17.52 20.60 2.72
+Gemini-1.5-Flash 17.9 17.50 10.77
+Gemini-1.5-Pro 18.25 18.20 12.69
+GPT-4o 19.20 17.12 16.19
+Gemini-1.5-Flash (CoT) 19.84 19.13 10.02
+Claude-3.5-Sonnet 22.26 25.33 9.21
+Multi-Turn Approaches (Gemini-1.5-Flash)
+Sk2code (N=2)∗∗ 19.41 20.45 11.81
+Self-Refine (N=2) 19.51 19.35 10.71
+BON (N=2) 21.45 20.1 13.5
+IAD (N=2) 24.78 23.01 15.29
+IAD-fb (N=2, K=2)∗∗ 24.86 23.4 14.69
+Self-Refine (N=4) 19.97 19.11 11.74
+Sk2code (N=4)∗∗ 20.41 21.46 12.67
+BON (N=4) 24.02 22.59 15.91
+IAD (N=4) 25.97 24.13 16.98
+IAD-fb (N=4, K=4)∗∗ 26.61 24.62 17.36
+Self-Refine (N=6) 19.89 18.91 11.61
+Sk2code (N=6)∗∗ 21.43 21.53 13.78
+BON (N=6) 25.75 22.91 17.67
+IAD (N=6) 26.75 24.91 19.12
+IAD-fb (N=6, K=6)∗∗ 27.95 24.99 19.01
+Table 3: Sketch2Code: Performance comparison between single-turn and multi-response generation approaches.
+For each of the multi-response generation method Layout score acts as the reference metric (temperature =0.6).
+Table demonstrate that IAD (Ours) consistently outperform SoTA baseline by >3-4% margin (absolute).
+∗∗ indicates those approaches leverage LLM-judge (oracle) feedback, although the LLM judges are similar but not exactly the same.
+Numbers taken from (Li et al., 2024b).
+N denotes the number of LLM calls for generating the HTML (>2000 tokens) and K represents the calls to LLM judge for getting feedback (<200 tokens).
+We have kept the judge-prompt similar to Sk2code (Li et al., 2024b)
 B.2 Sketch2code
@@
 Hence, we use Layout similarity as a verifier along with LLM-as-judge (Li et al., 2024b) to guide the generations for both BON (Beirami et al., 2024) and IAD.
-We report comparison with baseline single-turn approaches including SotA models GPT-4o, Claude-3, InternVL2, Gemini-1.5-Flash, CoT and variants along with multi-response generation approaches including BON, Sk2code and IAD (Ours).
-Single turn approaches even from SoTA models fail to match the layout structure, position of blocks, textual content, size of the blocks etc in the given user-sketch, causing a mismatch w.r.t to the reference layout as can be clearly seen in Figure and achieves a low score in-terms of all the three metrics in-comparison with multi-response generation approaches even with N=2.
-Best-of-N sampling (BON) with a weaker model Gemini-1.5-Flash improves over single-turn approaches and , with $N=4$ generations, it outperforms SoTA models with single-turn responses by a margin of 15-18%, by correctly identifying the block position, title block, overall layout structures etc.
+We report a comparison with baseline single-turn approaches, including SotA models GPT-4o, Claude-3, InternVL2, Gemini-1.5-Flash, CoT, and variants, along with multi-response generation approaches, including BON, Sk2code, and IAD (Ours).
+Single turn approaches even from SoTA models fail to match the layout structure, position of blocks, textual content, size of the blocks etc in the given user-sketch, causing a mismatch w.r.t to the reference layout as can be clearly seen in Figure and achieves a low score in-terms of all the three metrics in-comparison with multi-response generation approaches, even with N=2.
+Best-of-N sampling (BON) with a weaker model Gemini-1.5-Flash improves over single-turn approaches and, with $N=4$ generations, it outperforms SoTA models with single-turn responses by a margin of 15-18%, by correctly identifying the block position, title block, overall layout structure, etc.
 We see monotonic improvement in performance over the number of responses as the layout score improves from $20.41$ to $25.7$ with 6 responses.
@@
 Our proposed approach IAD, mitigates this gap by iteratively improving the responses and as shown in Table 1, it achieves a major improvement of $\leavevmode\nobreak\ 3-5\%$ (absolute improvement) from BON as well as single-turn SoTA Claude with just 2 iterations (eq : N=2) even with simpler model Gemini-1.5-Flash.
-At each iteration, we pass the best and worst HTML as a context along with instruction, for generating the next iteration.
+At each iteration, we pass the best and worst HTML as a context along with instructions for generating the next iteration.
 We observe IAD is able to learn fine-grained layout components, image semantics over iterations with the context of the Best and Worst HTML.
-We see that with increased iterations, performance of IAD improves reaching to a very high layout score of 26.75, outperforming all baselines with same generations.
+We see that with increased iterations, the performance of IAD improves, reaching a very high layout score of 26.75, outperforming all baselines with the same generations.
 We also report the Image and Text IoU scores while optimizing with the layout-score, to check for reward-overoptimization of the metric.